### v1

# On distinguishing coloring and some variants of proper coloring of graphs derived from subdivision operations

Abstract.
Let $G$ be a simple, finite, connected, and undirected graph, and $T$ be a finite tree. The middle graph $M(G)$ of $G$ is obtained from the subdivision graph $S(G)$ after joining pairs of subdivided vertices that lie on adjacent edges of $G$ and the central graph $C(G)$ of $G$ is obtained from $S(G)$ after joining all non-adjacent vertices of $G$. We show that if the order of $G$ is at least $4$, then $Aut(G)\cong Aut(C(G))\cong Aut(M(G))$(as abstract groups) and apply these results to obtain new sharp upper bounds of the distinguishing number and the distinguishing index of $C(G)$ and $M(G)$ inspired by an algorithm due to Kalinowski, Pilśniak, and Woźniak from 2016. Furthermore, we study the total distinguishing chromatic number of $C(G)$ and $S(G)$, use Latin squares to verify the AVD-total coloring conjecture for central graphs of regular graphs and some other classes of graphs (which is a partial progress towards answering an open question of Panda, Verma, and Keerti), and obtain new bounds of the total dominator chromatic number of $C(G)$ and $C(T)$.
Key words and phrases:
Automorphism groups, distinguishing coloring, total dominator coloring, adjacent vertex distinguishing total coloring, central graph, subdivision graph
2020 Mathematics Subject Classification:
Primary 05C15; Secondary 05C25, 05C76.
Corresponding author.
1. Introduction
Let $Aut(G)$ be the automorphism group of a graph $G=(V(G),E(G))$, $id_{G}$ be the trivial automorphism of $G$, and $f$ be a vertex coloring or an edge coloring or a total coloring (that is, a coloring of its vertices and edges) of $G$. An automorphism $\phi$ of $G$ preserves $f$ if the following holds: (1) $f(x)=f(\phi(x))$ for all $x\in V(G)$ if $f$ is a vertex coloring, (2) $f(e)=f(\phi(e))$ for all $e\in E(G)$ if $f$ is an edge coloring, (3) $f(x)=f(\phi(x))$ for all $x\in V(G)\cup E(G)$ if $f$ is a total coloring.
1.1. Distinguishing number and distinguishing index
Albertson and Collins [1] introduced the distinguishing number $D(G)$ of a graph $G$ as the least integer $d$ such that $G$ admits a vertex coloring with $d$ colors that is only preserved by $id_{G}$ and Kalinowski and Pilśniak [16] introduced the distinguishing index $D^{\prime}(G)$ of $G$ as the least integer $d$ such that $G$ admits an edge coloring with $d$ colors that is only preserved by $id_{G}$. In [4, Theorems 3.8 and 4.3], Alikhani and Soltani proved that if $G$ is a connected graph of order $n\geq 3$, then $D(S(G))\leq\lceil\sqrt{\Delta(G)}\rceil$ and $D^{\prime}(S(G))\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$ where $\Delta(G)$ is the maximum degree of $G$. Hamada and Yoshimura [13] and Vernold [31] introduced the concept of middle graphs and central graphs, respectively; classes of graphs derived from subdivision graphs. Since then, the notion of central graphs and middle graphs has been intensively studied in graph theory; see, for example, [5, 9, 11, 12, 18, 19, 23, 28].
Results: Let $G$ be a connected graph of order $n$. The following holds:
  1. (1)
    If $n\geq 3$, then $D(M(G))\leq\Delta(G)$,
  2. (2)
    $D^{\prime}(M(G))\leq 3$ for any positive integer $n$, and
  3. (3)
    If $n\geq 4$, then $D^{\prime}(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$, and $D(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$.
Moreover, the bounds mentioned in (1) and (3) are sharp (see Theorems 3.2, 3.4, 3.6).
1.2. Total distinguishing chromatic number
A proper total coloring of a graph $G$ is a total coloring of $G$ so that no adjacent vertices, no adjacent edges, and no incident vertices and edges get the same color. The least number of colors needed for a proper total coloring of $G$ is called its total chromatic number, denoted by $\chi^{\prime\prime}(G)$.
Behzad [6] and Vizing [33] independently posed the following conjecture.
Conjecture 1.1.
(Total-Coloring Conjecture (TCC)). For any graph $G$, $\chi^{\prime\prime}(G)\leq\Delta(G)+2$.
Recently, Panda et al. [28, Theorem 3.1] proved that TCC holds for the central graph of any graph of order $n\geq 5$. We observe an alternative short proof of the result using the following lemma.
Lemma 1.2.
([36, Lemma 2.4]) If $G$ is a graph of order $n$ that contains an independent set $S$ such that $|S|\leq n-\Delta(G)-1$, then $G$ satisfies TCC.
Theorem 1.3.
(Panda–Verma–Keerti;[28]) TCC holds for the central graph of any graph.
Short proof of Theorem 1.3.
Let $C(G)$ be the central graph of a graph $G=(V(G),E(G))$ where $V_{1}=V(G)$, $V_{2}=V(C(G))\backslash V(G)$, $|V_{1}|=n$, and $|V_{2}|=r$. If $n\geq 3$, then the order of $C(G)$ is $n+r$, $\Delta(C(G))=n-1$, and $C(G)$ contains an independent set $V_{2}$ of size $r=(n+r)-\Delta(C(G))-1$. By Lemma 1.2, $C(G)$ satisfies TCC. If $n\leq 2$, then it is trivial to verify. ∎
We say $G$ is of Type $j$ if $\chi^{\prime\prime}(G)=\Delta(G)+j$ for $j\in\{1,2\}$. In [15], Kalinowski, Pilśniak, and Woźniak introduced the total distinguishing chromatic number $\chi^{\prime\prime}_{D}(G)$ of a graph $G$ as the least integer $d$ such that $G$ has a proper total coloring with $d$ colors that is preserved only by $id_{G}$. We study the total distinguishing chromatic number of central and subdivision graphs.
Results: Let $G$ be a connected graph of order $n$. The following holds:
  1. (1)
    (Proposition 4.4) $\Delta(C(G))+1\leq\chi^{\prime\prime}(C(G))\leq\chi^{\prime\prime}_{D}(C(G))\leq\Delta(C(G))+2$.
  2. (2)
    (Theorem 4.5(1)) If $G$ is regular and $n\geq 5$ is odd, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
  3. (3)
    (Theorem 4.5(2)) If $G$ is regular and a non complete graph, $n\geq 5$ is even, and $\Delta(G)\leq\frac{2n-29}{6}$, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
  4. (4)
    (Proposition 4.9) $\chi^{\prime\prime}_{D}(S(G))$ is either $\Delta(S(G))+1$ or $\Delta(S(G))+2$.
1.3. AVD-total chromatic number
An adjacent vertex distinguishing (AVD) total coloring of a graph $G$ with edge set $E(G)$ is a proper total coloring $f$ such that for any edge $\{u,v\}\in E(G)$, we have $C_{G}(u)\neq C_{G}(v)$, where $C_{G}(u)=\{f(u)\}\cup\{f(\{u,v\}):\{u,v\}\in E(G)\}$ is the color class of $u$. The adjacent vertex distinguishing (AVD)-total chromatic number $\chi^{\prime\prime}_{a}(G)$ of $G$ is the minimum integer $k$ such that there exists an AVD-total coloring of $G$ using $k$ colors. Zhang et al. [37] introduced AVD-total-colorings and posed the following conjecture.
Conjecture 1.4.
(AVD-TCC). For any graph $G$, $\chi^{\prime\prime}_{a}(G)\leq\Delta(G)+3$.
We say $G$ is of AVD-Type $i$ if $\chi^{\prime\prime}_{a}(G)=\Delta(G)+i$ for $i\in\{1,2,3\}$. The Conjecture 1.4 for general graphs $G$ with $\Delta(G)=3$ has been verified by Wang [34] and Chen [7]. Lu et al. [22] validated the conjecture for all graphs $G$ with $\Delta(G)=4$. The Conjecture 1.4 is also known to be true for many families of graphs such as 4-regular graphs [26], hypercubes [8], complete equipartite graphs [24], indifference graphs [27], split graphs [32], and several other classes of graphs. Panda et al. [28] obtained the AVD-total chromatic number of central graphs for some known families of graphs and mentioned the following open question (see [28, Section 5]).
Question 1.5.
(Panda et al. [28]) Does AVD-TCC hold for the central graph of any graph?
In [28, Section 5], it was mentioned that it would be interesting to solve the AVD-total coloring classification problem on the central graph of any graph. We have made partial progress in this direction.
Results: The following holds:
  1. (1)
    (Theorem 5.1) If $G$ is a regular graph of even order $n\geq 5$, then $C(G)$ is AVD-Type 2 and if $G$ is a regular graph of order $n\geq 5$, then $C(G)$ satisfies AVD-TCC.
  2. (2)
    (Theorem 5.3) If $\Delta(G)\geq 5$, then the subdivision graph $S(G)$ of $G$ is AVD-Type 1.
  3. (3)
    (Proposition 5.5 (4)) If two graphs $G_{1}$ and $G_{2}$ have the same order and $G_{1}+G_{2}$ is the join of $G_{1}$ and $G_{2}$, then $C(G_{1}+G_{2})$ satisfies AVD-TCC.
1.4. Total dominator chromatic number
The notion of total domination number was introduced by Cockayne, Dawes and Hedetniemi in [10] and has been studied for central graphs and middle graphs, see [9, 18, 19, 20, 21]. In particular, Kazemnejad and Moradi [19] and Chen et al. [9] studied sharp upper bounds for the total domination number of central graphs of connected graphs and trees. Motivated by the relation between the chromatic number and the total domination number, the notion of the total dominator chromatic number was studied in [17]. A total dominator coloring, briefly TDC, of a graph $G$ is a proper vertex coloring of $G$ in which each vertex of $G$ is adjacent to every vertex of some color class. The total dominator chromatic number $\chi^{t}_{d}(G)$ of $G$ is the minimum number of color classes in a TDC of $G$. The total dominator chromatic number was studied for middle graphs by Kazemnejad et al. [20], $k$-subdivision of graphs by Alikhani et al. [2], and central graphs by Kazemnejad et al. [18]. In [18, Theorem 2.1], Kazemnejad and Kazemi proved that the following bounds are sharp if $G$ is a connected graph of order $n\geq 2$:
  • $\lfloor 2n/3\rfloor+1\leq\chi^{t}_{d}(C(G))\leq\Delta(C(G))+2$ if $\Delta(G)\leq n-2$.
  • $\chi^{t}_{d}(C(G))\leq n+\lceil n/2\rceil$ if $G$ has a Hamiltonian path.
Results: Let $G$ be a connected graph of order $n\geq 5$. The following holds:
  1. (1)
    If $\Delta(G)\leq n-3$ and $\overline{G}$ is the complement of $G$, then $\chi^{t}_{d}(\overline{G})\leq\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
  2. (2)
    If $G$ is a tree, then $\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
Moreover, the upper bounds of $\chi^{t}_{d}(C(G))$ mentioned in (1) and (2) are sharp (see Theorem 6.1).
1.5. Tables of results
Let $C_{n}$ be a cycle graph of $n$ vertices, $P_{n}$ be a path of $n$ vertices, $K_{n}$ be a complete graph of $n$ vertices, and $K_{n,m}$ be a complete bipartite graph with bipartitions of sizes $n$ and $m$. A claw is the graph $K_{1,3}$. The degree of a vertex $v\in V(G)$ in $G=(V(G),E(G))$, denoted by $deg_{G}(v)$, is the number of edges that emerge from $v$. We say $G$ is a regular graph if all vertices of $G$ have the same degree. The degree of a regular graph $G$, denoted by $d(G)$, is the same as the maximum degree $\Delta(G)$ and the minimum degree $\delta(G)$. Let $N_{G}(v)=\{u\in V(G):\{u,v\}\in E(G)\}$ for any $v\in V(G)$.
Graph $D^{\prime}(H)$ $D(H)$
$H=C(G),n\geq 4$ $\leq\lceil\sqrt{\Delta(G)}\rceil$ (Thm. 3.2) $\leq\lceil\sqrt{\Delta(G)}\rceil$ (Thm. 3.4)
$H=M(G),n\geq 3$ $\leq 3$ (Thm. 3.6(2)) $\leq\Delta(G)$ (Thm. 3.6(1))
$H=S(G)$ $\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$ ([4, Thm. 4.3]) $\leq\lceil\sqrt{\Delta(G)}\rceil$ ([4, Thm. 3.8])
Graph $\chi^{\prime\prime}_{a}(H)$ $\chi^{\prime\prime}_{D}(H)$
$H=S(G),\Delta(G)\geq 5$ $\Delta(H)+1$ (Thm. 5.3) $\leq\Delta(H)+2$ (Prop. 4.9)
$H=C(G),n\equiv 0$(mod 2), $G\not\cong K_{n}\text{ is regular}$, $n\geq 5$, and $\Delta(G)\leq\frac{2n-29}{6}$ $\Delta(H)+2$ (Thm. 5.1) $\Delta(H)+1$ (Thm. 4.5)
[…]
$H=C(K_{m,r})$ $\leq\Delta(H)+3$ (Prop. 5.5(2)) $\leq\Delta(H)+2$ (Prop. 4.4)
$H=C(G_{1}+G_{2})$, $G_{1},G_{2}\in\{P_{n},C_{n},K_{n},K_{n,r}\}$ $\leq\Delta(H)+3$ (Prop. 5.5(3)) $\leq\Delta(H)+2$ (Prop. 4.4)
$H=C(H_{1}+H_{2})$, order of $H_{1}$ = order of $H_{2}$ $\leq\Delta(H)+3$ (Prop. 5.5(4)) $\leq\Delta(H)+2$ (Prop. 4.4)
Graph $\chi^{t}_{d}(H)$
$H=C(G),n\geq 5,\Delta(G)\leq n-3$ $\geq\chi^{t}_{d}(\overline{G})$, $\leq\Delta(H)+1$ (Thm. 6.1(1))
$H=C(T),T\text{ is a tree, }\text{order of }T\geq 5$ $\leq\Delta(H)+1$ (Thm. 6.1(2))
$H=M(G),n\geq 2,G\text{ has $m$ edges}$ $\leq n+m-1$ ([20, Thm. 2.3])
$H=S(G),G\text{ has $m$ edges}$ $\leq 2m$ ([2, Thm. 2.2])
$H=C(G),n\geq 3,\Delta(G)\leq n-2$ $\leq\Delta(H)+2$ ([18, Thm. 2.1])
$H=C(G),n\geq 2,\text{ $G$ has a Hamiltonian path}$ $\leq n+\lceil n/2\rceil$ ([18, Thm. 2.1])
In the above tables, we list the results obtained in this manuscript, where $n$ denotes the order of $G$.
2. Automorphism groups
Definition 2.1.
The subdivision graph $S(G)$ of a graph $G=(V(G),E(G))$ is a bipartite graph with the set of vertices $V(S(G))=V_{1}\cup V_{2}$ where $V_{2}=\{w_{v,v^{\prime}}:w_{v,v^{\prime}}\text{ subdivide the edge }\{v,v^{\prime}\}\in E(G)\}$ and $V_{1}=V(G)$, and the set of edges $E(S(G))=\{\{v,w_{v,v^{\prime}}\},\{w_{v,v^{\prime}},v^{\prime}\}\}:\{v,v^{\prime}\}\in E(G)\}$.
Definition 2.2.
The central graph $C(G)$ of a graph $G=(V(G),E(G))$ of order $n$ and size $m$ is a graph of order $n+m$ and size ($n\choose{2}$ + $m$ that is obtained by subdividing each edge of $G$ exactly once and joining all the non-adjacent vertices of $G$ in $C(G)$. In particular, $C(G)$ is a graph with the set of vertices $V(C(G))=V(S(G))$ and the set of edges $E(C(G))=E(\overline{G})\cup\{\{v,w_{v,v^{\prime}}\},\{w_{v,v^{\prime}},v^{\prime}\}\}:\{v,v^{\prime}\}\in E(G)\}$ where $\overline{G}$ is the complement of $G$.
In this section, we sketch some lemmas concerning $Aut(C(G))$ and $Aut(M(G))$ where $M(G)$ is the middle graph of $G$. Applying these lemmas, we will prove Theorems 3.2, 3.4, and 4.5.
Lemma 2.3.
If $\psi\in Aut(C(G))$ such that $\psi(u)=u$ for all $u\in V_{1}$, then $\psi=id_{C(G)}$.
Proof.
First, we note that $N_{C(G)}(a)\neq N_{C(G)}(b)$ for any $a,b\in V_{2}$. Let $x\in V_{2}$. If $\psi(x)=y$ for some $y\in V_{2}\backslash\{x\}$, then $\psi(N_{C(G)}(x))=N_{C(G)}(y)$ since automorphisms preserve adjacency. Since for every $v\in V_{1}$, $\psi(v)=v$, we have $\psi(N_{C(G)}(x))=N_{C(G)}(x)$. Thus, $N_{C(G)}(x)=N_{C(G)}(y)$, which is a contradiction. ∎
Lemma 2.4.
Let $G$ be a connected graph of order $n\geq 4$. If $\varphi\in Aut(C(G))$, then $\varphi(V_{1})=V_{1}$ and $\varphi(V_{2})=V_{2}$.
Proof.
In $C(G)$, each vertices of $V_{1}$ is of degree $n-1\geq 3$ and each vertices of $V_{2}$ is of degree $2$. The rest follows from the fact that $deg_{C(G)}(\phi(v))=deg_{C(G)}(v)$ for any $v\in V(C(G))$. ∎
Lemma 2.5.
Let $G$ be a connected graph of order $n\geq 4$. Then $Aut(G)\cong Aut(C(G))$.
Proof.
Let $\alpha\in Aut(G)$. Define the mapping $\pi_{\alpha}:V(C(G))\rightarrow V(C(G))$ as follows:
$\pi_{\alpha}(w)=\begin{cases}\alpha(w)&\text{if }w\in V(G)\\ w_{\alpha(x),\alpha(y)}&\text{if }w\in V(C(G))\backslash V(G),w=w_{x,y},x,y\in V(G).\end{cases}$
We can see that $\pi_{\alpha}\in Aut(C(G))$. If $\{x,y\}\in E(C(G))$ then $\{x,y\}\in E(\overline{G})$ or $\{x,y\}\in E(S(G))$. By the definition of $\pi_{\alpha}$, if $\{x,y\}\in E(S(G))$ then $\{\pi_{\alpha}(x),\pi_{\alpha}(y)\}\in E(S(G))$. Moreover, we have $\{x,y\}\in E(\overline{G})\iff\{x,y\}\not\in E(G)\iff\{\alpha(x),\alpha(y)\}\not\in E(G)\iff\{\alpha(x),\alpha(y)\}\in E(\overline{G})\iff\{\pi_{\alpha}(x),\pi_{\alpha}(y)\}\in E(\overline{G})$. Thus, if $H=\{\pi_{\alpha}:\alpha\in Aut(G)\}$, then $H\subseteq Aut(C(G))$. We show that $Aut(C(G))\subseteq H$. Let $f\in Aut(C(G))$. By Lemma 2.4, we have $f(V_{1})=V_{1}$ since $G$ is a connected graph of order at least $4$. If $\alpha=f\restriction V_{1}$, then $\alpha\in Aut(G)$. Let $\pi_{\alpha}$ be the induced automorphism of $\alpha$ as defined above. If we let $l=f^{-1}\pi_{\alpha}$, then $l$ is an automorphism of $C(G)$ such that $l(v)=v$ for every $v\in V_{1}$. By Lemma 2.3, $l=id_{C(G)}$. Consequently, $f=\pi_{\alpha}$. Therefore, $Aut(C(G))\subseteq H$, and so $H=Aut(C(G))$. It is trivial to see that $H\cong Aut(G)$, so we are done. ∎
The following definition is due to Hamada and Yoshimura [13].
Definition 2.6.
Let $G=(V(G),E(G))$ be a graph where $V(G)=\{v_{1},...,v_{p}\}$ for some integer $p\geq 1$. To $G$, we add $p$ new vertices $\{u_{1},...,u_{p}\}$ and $p$ new edges $\{u_{i},v_{i}\}$. The new graph is the endline graph of $G$ denoted by $G^{+}$. The edges $\{u_{i},v_{i}\}$ are the endline edges of $G^{+}$. Let $U_{1}=V(G)$ and $U_{2}=V(G^{+})\backslash V(G)$. Let $L(G)$ be the line graph of $G$.
Fact 2.7.
The following holds:
  1. (1)
    (Hamada–Yoshimura; [13]) Let $G$ be any graph. Then $M(G)\cong L(G^{+})$.
  2. (2)
    (Sabidussi; [30]) If $G$ is a connected graph where $G\not\in\{P_{2},Q,L(Q)\}$ (see Fig. 1). Then $Aut(G)\cong Aut(L(G))$ where $L(G)$ is the line graph of $G$.
  3. (3)
    (Mirafzal; [25, Theorem 4]) If $G\not\cong C_{n}$ for any $n\geq 3$, then $Aut(S(G))\cong Aut(G)$.
[…]
Figure 1. Graphs $L(Q),Q,$ and $C_{6}^{+}$.
Lemma 2.8.
If $\psi\in Aut(G^{+})$ such that $\psi(u)=u$ for all $u\in U_{1}$, then $\psi=id_{G^{+}}$.
Proof.
Let $u_{i}\in U_{2}$. If $\psi(u_{i})=u_{j}$ for some $u_{j}\in U_{2}\backslash\{u_{i}\}$, then $\psi(v_{i})=v_{j}$ since the automorphisms preserve adjacency, which is a contradiction to the given assumption. ∎
Lemma 2.9.
If $G$ is a non-trivial connected graph and $\varphi\in Aut(G^{+})$, then $\varphi(U_{i})=U_{i}$ for $i\in\{1,2\}$.
Proof.
In $G^{+}$, each vertex of $U_{1}$ has a degree at least $2$ and each vertex of $U_{2}$ has a degree $1$. ∎
Lemma 2.10.
Let $G$ be a connected graph of order $n\geq 3$. Then $Aut(G)\cong Aut(M(G))$.
Proof.
Since $G^{+}\not\in\{Q,L(Q),P_{2}\}$ as $G^{+}$ has $n$ vertices of degree one (see Fig. 1), $Aut(M(G))\cong Aut(L(G^{+}))\cong Aut(G^{+})$ by Fact 2.7. Let $\alpha\in Aut(G)$. Define the mapping $\pi_{\alpha}:V(G^{+})\rightarrow V(G^{+})$ by the following rule:
$\pi_{\alpha}(w)=\begin{cases}\alpha(v_{i})&\text{if }w=v_{i}\in U_{1}\\ u_{j}&\text{if }w=u_{i}\in U_{2},v_{j}=\alpha(v_{i}).\end{cases}$
Then $\pi_{\alpha}\in Aut(G^{+})$, where $\pi_{\alpha}(U_{i})=U_{i}$ for $i\in\{1,2\}$. The rest follows by applying Lemmas 2.8, 2.9, and the arguments in the proof of Lemma 2.5. ∎
By Lemmas 2.5, 2.10, and Fact 2.7(2,3), we have the following result.
Theorem 2.11.
If $G$ is a connected graph of order $n\geq 5$ and $G\not\cong C_{n}$, then $Aut(G)\cong Aut(L(G))\cong Aut(S(G))\cong Aut(C(G))\cong Aut(M(G))$.
3. Distinguishing number and distinguishing index
In [15], Kalinowski, Pilśniak, and Woźniak introduced the total distinguishing number $D^{\prime\prime}(G)$ of a graph $G$ as the least integer $d$ such that $G$ has a total coloring (not necessarily proper) with $d$ colors that is preserved only by $id_{G}$.
Fact 3.1.
Let $G$ be a connected graph of order $n$. The following holds:
  1. (1)
    (Kalinowski, Pilśniak, and Woźniak; [15, Theorem 2.2]) If $n\geq 3$, then $D^{\prime\prime}(G)\leq\lceil\sqrt{\Delta(G)}\rceil$.
  2. (2)
    (Alikhani and Soltani; [4, Theorems 3.8 and 4.3]) If $n\geq 3$, then $D(S(G))\leq\lceil\sqrt{\Delta(G)}\rceil$ and $D^{\prime}(S(G))\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$.
  3. (3)
    (Kalinowski and Pilśniak; [16, Propositions 5 and 6]) $D^{\prime}(C_{t})=2$ and $D^{\prime}(K_{t})=2$ for $t\geq 6$.
  4. (4)
    (Alikhani and Soltani; [3, Theorem 2.3]) If $G\not\in\{P_{2},L(Q)\}$, then $D(L(G))=D^{\prime}(G)$ (see Fig. 1).
  5. (5)
    (Kalinowski and Pilśniak; [16]) If $n\geq 3$, then $D^{\prime}(G)\leq\Delta(G)$ unless $G$ is $C_{3},C_{4},$ or $C_{5}$.
Theorem 3.2.
Let $G$ be a connected graph of order $n\geq 4$. Then $D^{\prime}(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$.
Proof.
Since $G$ is connected, we have $\Delta(G)\geq 2$. Hence, we have at least $\lceil\sqrt{2}\rceil=2$ colors available to color $C(G)$. For $k\geq 0$, let $S_{k}(x)$ denote a sphere of radius $k$ with a center $x$ in $G$, that is, the set of all vertices at distance $k$ from $x$.
Case (i): $G=K_{n}$ or $C_{n}$.
If $n\geq 6$, then $D^{\prime}(G)=2\leq\lceil\sqrt{\Delta(G)}\rceil$ by Fact 3.1(3). We assume $n$ is either $4$ or $5$. In both cases, we can color the edges of $C(G)$ with two available colors so that the edge coloring is preserved only by $id_{G}$ (see Fig. 2).
$C(C_{4})$$C(C_{5})$$C(K_{4})$$C(K_{5})$
Figure 2. Edge colorings of $G$ with $2$ colors that are only preserved by $id_{G}$ where $G\in\{C(C_{4}),C(C_{5}),C(K_{4}),C(K_{5})\}$.
Case (ii): $G$ contains a cycle but $G\not\in\{C_{n},K_{n}\}$.
Thus, we can choose a vertex $v_{0}$ lying on a cycle such that the sphere $S_{2}(v_{0})$ is non-empty. Consider a BFS-spanning tree $T$ of $G$ rooted at $v_{0}$. For a vertex $v\in T$, we denote
$X_{v,u}=(\{v,w_{v,u}\},\{w_{v,u},u\})$ if $w_{v,u}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)$
and
$N(v)=\{X_{v,u}:w_{v,u}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)\}$.
Step 1: We color all edges of the subdivision graph of $T$ and certain edges of $C(G)$.
  • We pick $v_{1},v_{2}\in S_{1}(v_{0})$ and assign the pair $(1,1)$ of colors to $X_{v_{0},v_{1}}$ and $X_{v_{0},v_{2}}$. Since the degree of $v_{0}$ is at most $\Delta(G)$ and the pair $(1,1)$ is used twice to color the elements of $N(v_{0})$, we can color every pair of $N(v_{0})\setminus\{X_{v_{0},v_{1}},X_{v_{0},v_{2}}\}$ with different pairs of colors from $(1,1)$. We modify an algorithm due to Kalinowski–Pilśniak–Woźniak [15] to color all the edges of the subdivision graph of $T$ so that $v_{0}$ is the only vertex such that the pair $(1,1)$ appears twice in $N(v_{0})$. Hence $v_{0}$ will be fixed by every color-preserving automorphism, and all vertices in $N(v_{0})$ will be fixed, except possibly $v_{1}$ and $v_{2}$. For $i\in\{1,2\}$, let
    $T_{v_{i}}=\{X_{v_{i},u}\in N(v_{i}):u\in S_{2}(v_{0})\}.$
    In order to distinguish $v_{1}$ and $v_{2}$, we need to take the following steps:
    • ($\ast$)
      We color the elements of $T_{v_{1}}$ and $T_{v_{2}}$ with the same pairs so that none of them contains the pair $(1,1)$.
[…]
      If $X_{v_{1},u}\in T_{v_{1}}$, then $u\in S_{2}(v_{0})$. Thus, $\{v_{0},u\}\in E(\overline{G})\subseteq E(C(G))$. We color all such edges with 1. On the other hand, we color all edges of the set $\{\{v_{0},u\}\in E(C(G)):X_{v_{2},u}\in T_{v_{2}}\}$ with 00.
Claim 3.3.
If $f$ is an edge coloring of $C(G)$ and $v_{0}$ is fixed by a $f$-preserving automorphism $\phi$, then $\phi(v_{i})=v_{i}$ for $i\in\{1,2\}$.
Proof.
[…]
  • Suppose that for some integer $k$, every element $u\in\bigcup_{i\leq k}S_{i}(v_{0})$ is colored so that it is fixed by every color-preserving automorphism. For every $u\in S_{k}(v_{0})$, color $X_{u,z}$, where $z$ is a child of $u$ in $T$, with distinct color pairs except $(1,1)$. Hence, all neighbors in $N(u)\cap S_{k+1}(v_{0})$ will be fixed. So, the elements of $\bigcup_{i\leq k+1}S_{i}(v_{0})$ are also fixed by every color-preserving automorphism. Continuing in this fashion, we color all the edges of $S(T)$.
Step 2: We color all the remaining edges in $C(G)$ with 00.
Step 3: Using Lemmas 2.3 and 2.4, we prove that the coloring that we defined above (say $c$) is preserved only by $id_{C(G)}$.
  • Let $\psi\in Aut(C(G))$ so that $\psi$ preserves $c$. By Lemma 2.4, $\psi(V_{i})=V_{i}$ for $i\in\{1,2\}$. Since all vertices of $T$ are fixed by $\psi$, we have $\psi(u)=u$ for all $u\in V_{1}$. By Lemma 2.3, $\psi$ is $id_{C(G)}$.
Case (iii): $G$ is a tree.
Then $G$ has a central vertex or a central edge that is fixed by every automorphism of $G$.
  1. Case A: If $G$ has a central vertex $v_{0}$, then every element of $N(v_{0})$ obtains a different pair of colors (where $N(v)$ is defined in $Case(ii)$ for any vertex $v$ in the tree $G$). This is possible as we have at least $\Delta(G)$ different pairs of colors available. Hence, all vertices adjacent to $v_{0}$ in $G$ are fixed by every automorphism of $G$, preserving this coloring. If $v\in S_{k}(v_{0})$, then all the elements of $N(v)\backslash\{X_{v,u}:u\in S_{k-1}(v_{0})\}$ obtain a different pair of colors. By induction on the distance from $v_{0}$, all vertices of $G$ are fixed.
  2. Case B: If $G$ has a central edge $e_{0}=\{x,y\}$, let $T_{1}$, $T_{2}$ be subtrees obtained by deleting $e_{0}$. If we put different colors on the edges $\{x,w_{x,y}\}$ and $\{w_{x,y},y\}$, then $x$ and $y$ are fixed by every automorphism. We color $T_{i}$ using the same method as in Case A for $i\in\{1,2\}$.
Theorem 3.4.
Let $G$ be a connected graph of order $n\geq 4$. Then $D(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$.
Proof.
In view of Fact 3.1(1), it suffices to show that $D(C(G))\leq D^{\prime\prime}(G)$. Let $f:V(G)\cup E(G)\to\{1,...,D^{\prime\prime}(G)\}$ be a total coloring of $G$ (not necessarily proper) that is preserved only by $id_{G}$. We define $\tilde{f}:V(C(G))\to\{1,...,D^{\prime\prime}(G)\}$ so that $\tilde{f}(x)=f(x)$ for all $x\in V_{1}$ and $\tilde{f}(w_{x,y})=f(\{x,y\})$ for all $w_{x,y}\in V_{2}$. We claim that $\tilde{f}$ is preserved only by $id_{C(G)}$. Let $\psi$ be a $\tilde{f}$-preserving automorphism of $C(G)$. If $H=\{\pi_{\alpha}:\alpha\in Aut(G)\}$ where $\pi_{\alpha}$ is the induced automorphism of $\alpha\in Aut(G)$, which is defined in the proof of Lemma 2.5, then $H=Aut(C(G))\cong Aut(G)$ (see the proof of Lemma 2.5). Thus, $\psi=\pi_{\alpha}$ for some $\alpha\in Aut(G)$ where $\psi(V_{1})=V_{1}$ and $\psi(V_{2})=V_{2}$. Since $\psi$ is a $\tilde{f}$-preserving automorphism of $C(G)$, we can see that $\alpha$ preserves $f$ for the following reasons:
  1. (1)
    if $v\in V_{1}$, then $\tilde{f}(\psi(v))=\tilde{f}(v)\implies f(\psi(v))=f(v)\implies f(\alpha(v))=f(\pi_{\alpha}(v))=f(v)$.
  2. (2)
    if $w_{x,y}\in V_{2}$, then $\tilde{f}(\psi(w_{x,y}))=\tilde{f}(w_{x,y})\implies\tilde{f}(\pi_{\alpha}(w_{x,y}))=f(\{x,y\})\implies\tilde{f}(w_{\alpha(x),\alpha(y)})=f(\{x,y\})\implies f(\{\alpha(x),\alpha(y)\})=f(\{x,y\})\implies f(\alpha(\{x,y\}))=f(\{x,y\})$.
So, $\alpha=id_{G}$ since $f$ is preserved only by $id_{G}$. Consequently, $\psi=\pi_{\alpha}=id_{C(G)}$. ∎
[…]
Figure 3. A vertex coloring of $C(K_{1,4})$ with $2$ colors and an edge coloring of $C(K_{1,5})$ with $3$ colors that are only preserved by trivial automorphisms.
Remark 3.5.
For any integer $n\geq 1$, we have $D^{\prime}(C(K_{1,n}))=D(C(K_{1,n}))=\lceil\sqrt{n}\rceil=\lceil\sqrt{\Delta(K_{1,n})}\rceil$. This shows that the bounds mentioned in Theorems 3.2 and 3.4 are sharp (see Fig. 3).
Theorem 3.6.
Let $G$ be a connected graph of order $n$. The following holds:
  1. (1)
    If $n\geq 3$, then $D(M(G))\leq\Delta(G)$. Moreover, the bound is sharp.
  2. (2)
    $D^{\prime}(M(G))\leq 3$ for any positive integer $n$.
Proof.
(1). By Facts 3.1(4) and 2.7 (1), $D(M(G))=D(L(G^{+}))=D^{\prime}(G^{+})$.
Case(i): $G\not\cong C_{n}$ for any $n\geq 3$. We show that $D^{\prime}(G^{+})\leq D^{\prime}(G)$. Then $D^{\prime}(G^{+})\leq\Delta(G)$ by Fact 3.1(5). Let $g:E(G)\rightarrow\{1,...,D^{\prime}(G)\}$ be a coloring of $E(G)$ that is preserved only by $id_{G}$ and $E$ be the set of all endline edges of $G^{+}$. Define the mapping $g^{\prime}:E(G^{+})\rightarrow\{1,...,D^{\prime}(G)\}$ as follows:
$g^{\prime}(e)=\begin{cases}1&\text{if}\,e\in E,\\ g(e)&\text{if}\,e\not\in E.\end{cases}$
We show that $g^{\prime}$ is preserved only by $id_{G^{+}}$. If $\psi\in Aut(G^{+})$ preserves $g^{\prime}$, then $\psi$ maps edges from $E$ to $E$ by Lemma 2.9. So, $\psi\restriction G\in Aut(G)$ preserves $g$. Thus, $\psi\restriction G=id_{G}$. By Lemma 2.8, $\psi=id_{G^{+}}$.
Case(ii): $G\cong C_{n}$ for some $n\geq 3$. Then $D(M(G))=D^{\prime}(G^{+})=2=\Delta(G)$. This also shows that the bound mentioned in (1) is sharp.
(2). Since every line graph is claw-free, and the distinguishing index of a connected claw-free graph is at most $3$ (see [29, Theorem 3.4]), we have $D^{\prime}(M(G))=D^{\prime}(L(G^{+}))\leq 3$. ∎
4. Total distinguishing chromatic number
Definition 4.1.
A Latin square of order $k$ is a $k\times k$ array based on the elements $1,2,...,k$ such that each element occurs exactly once in each row and exactly once in each column. A Latin square $M=[m_{i,j}]$ of order $k$ is said to be commutative if $m_{i,j}=m_{j,i}$, for $1\leq i,j\leq k$ and $M$ is said to be idempotent if $m_{i,i}=i$, for $1\leq i\leq k$. If the rows of $M$ are just cyclic permutations (one shift of the elements to the left) of the previous row, then $M$ is said to be anti-circulant.
1 5 2 6 3 7 4
5 2 6 3 7 4 1
2 6 3 7 4 1 5
6 3 7 4 1 5 2
3 7 4 1 5 2 6
7 4 1 5 2 6 3
4 1 5 2 6 3 7
Table 1. An anti-circulant idempotent commutative Latin square of order $7$.
Definition 4.2.
A proper edge coloring of a graph $G=(V(G),E(G))$ is a coloring of $E(G)$ so that any two adjacent edges have different colors. The chromatic index of $G$, denoted by $\chi^{\prime}(G)$, is the least number of colors needed for a proper edge coloring of $G$. Fix an integer $k$. We say that $G$ is $k$-edge choosable if for any assignment $L={L(e)}_{e\in E(G)}$ of the lists of available colors to the edges of $G$, there is a proper edge coloring $f$ of $G$ so that $f(e)\in L(e)$ and $|L(e)|=k$ for all $e\in E(G)$. The list chromatic index of $G$, denoted by $\chi_{L}^{\prime}(G)$, is the minimum integer $k$ such that $G$ is $k$-edge choosable.
Fact 4.3.
The following holds:
  1. (1)
    (Galvin; [14]) If $G$ is a bipartite graph, then $\chi^{\prime}(G)=\chi_{L}^{\prime}(G)=\Delta(G)$.
  2. (2)
    An idempotent commutative Latin square (ICLS) of order $k$ exists if and only if $k$ is odd.
  3. (3)
    If $M(k)=[m_{i,j}]$ is a Latin square where $m_{i,j}\equiv(i+j)k$ (mod $12k-$ 1), $11\leq m_{i,j}\leq 2k-$ 1, for $11\leq i,j\leq 2k-$ 1, then $M(k)$ is an anti-circulant ICLS of order $12k-$ 1.
  4. (4)
    (Xie–He;[35, Theorem 1]) If $G$ is a regular graph of even order $n$ and $\Delta(G)\geq\frac{2n}{3}+\frac{23}{6}$, then $G$ has TCC.
  5. (5)
    (Kalinowski–Pilśniak–Woźniak;[15]) Every connected graph $G$ satisfies the inequality $\chi^{\prime\prime}_{D}(G)\leq\chi^{\prime\prime}(G)+1$. Furthermore, $\chi_{D}^{\prime\prime}(G)=\chi^{\prime\prime}(G)$ if $\chi^{\prime\prime}(G)\geq\Delta(G)+2$.
4.1. Central graphs
Proposition 4.4.
If $G$ is a connected graph, then $\chi^{\prime\prime}_{D}(C(G))$ is $\Delta(C(G))+1$ or $\Delta(C(G))+2$.
Proof.
It is evident that $\Delta(G)+1\leq\chi^{\prime\prime}(G)\leq\chi^{\prime\prime}_{D}(G)$ for any graph $G$. By Theorem 1.3 and Fact 4.3(5), we have $\chi^{\prime\prime}_{D}(C(G))\leq\Delta(C(G))+2$. ∎
Theorem 4.5.
Let $G$ be a connected regular graph of order $n\geq 5$. The following holds:
  1. (1)
    If $n$ is odd, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
  2. (2)
    If $n$ is even, $G\not\cong K_{n}$, and $\Delta(G)\leq\frac{2n-29}{6}$, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
Consequently, $C(G)$ is Type 1 if $G$ satisfies the requirement of (1) or (2).
Proof.
We note that $\Delta(C(G))=n-1$. We recall that $V_{1}=V(G)$ and $V_{2}=V(C(G))\backslash V(G)$.
[…]
Claim 4.6.
There is a total coloring $g$ of $C(G)$ using $\Delta(C(G))+1$ colors.
Proof.
Let $V_{1}=\{u_{1},...,u_{p}\}$ and $V_{2}=\{v_{1},...,v_{q}\}$. First, we define a proper total coloring $\pi$ of $\overline{G}$ from $M(k)$ with $\Delta(C(G))+1$ colors. Let $\pi(u_{i})=m_{i,i}$ for all $1\leq i\leq n$ and $\pi(\{u_{i},u_{j}\})=m_{i,j}$ for all $1\leq i,j\leq n,i=j$ be such that $\{u_{i},u_{j}\}\in E(\overline{G})$. Next, we use Fact 4.3(1) to define a proper edge coloring of the bipartite graph $B=C(G)\backslash E(\overline{G})$. We note that for each vertex $u_{i}\in V_{1}$, there are $\Delta(C(G))+1-(deg_{\overline{G}}(u_{i})+1)=\Delta(C(G))-deg_{\overline{G}}(u_{i})$ colors that are not used from $X(u_{i})=\{m_{i,j}:1\leq j\leq\Delta(C(G))+1\}$, i.e., the $i^{th}$- row of $M(k)$. Since $G$ is a regular graph, we have $\Delta(C(G))-deg_{\overline{G}}(u_{i})=\Delta(C(G))-\Delta(\overline{G})=\Delta(B)$. Fix a vertex $u_{i}\in V_{1}$ and an edge $\{u_{i},v_{j}\}\in E(B)$. Let $V(u_{i})=\{m_{i,j}:\{u_{i},u_{j}\}\in E(\overline{G}),1\leq j\leq\Delta(C(G))+1\}$ and $L(\{u_{i},v_{j}\})=X(u_{i})-V(u_{i})$. Assume that $L=L(\{u_{i},v_{j}\})_{\{u_{i},v_{j}\}\in E(B)}$ is an assignment of lists of available colors for the edges of the bipartite graph $B$. By Fact 4.3(1), there exists a proper edge coloring $\pi_{1}$ of $B$ such that $\pi_{1}(e)\in L(e)$ for all $e\in E_{B}$. We define a total coloring $g$ of $C(G)$ using $\Delta(C(G))+1$ colors as follows:
  • Let $g(\{u_{i},v_{j}\})=\pi_{1}(\{u_{i},v_{j}\})$ for all $\{u_{i},v_{j}\}\in E(B)$,
  • Let $g(v)=\pi(v)$ for all $v\in V_{1}$ and $g(e)=\pi(e)$ for all $e\in E(\overline{G})$,
  • If $v_{k}=w_{u_{i},u_{j}}\in V_{2}$ is a vertex that subdivides the edge $\{u_{i},u_{j}\}\in E(G)$, then since $n\geq 5$, let $g(v_{k})$ be any color from the set $\{m_{i,1},...,m_{i,n}\}\backslash\{g(\{v_{k},u_{i}\}),g(\{v_{k},u_{j}\}),g(u_{i}),g(u_{j})\}$.
We can see that the coloring $g$ is a proper total coloring of $C(G)$. ∎
Claim 4.7.
$g$ is only preserved by $id_{C(G)}$.
Proof.
Let $\phi\in Aut(C(G))$ be such that $\phi$ preserves $g$. By Lemma 2.4, we have $\phi(V_{1})=V_{1}$ and $\phi(V_{2})=V_{2}$. Define $f:=g\restriction(V(\overline{G})\cup E(\overline{G}))$ and $\psi:=\phi\restriction V_{1}$. Then $\phi\in Aut(\overline{G})$. Since $\phi$ preserves $g$, we have the following:
  • For all $v\in V_{1}$, $g(\phi(v))=g(v)$. Thus, $f(\phi(v))=g(\phi(v))=g(v)=f(v)$ since $\phi(v),v\in V_{1}$.
  • If $\{v,v^{\prime}\}\in E(\overline{G})$ where $v,v^{\prime}\in V_{1}$, then $\{\phi(v),\phi(v^{\prime})\}\in E(\overline{G})$ where $\phi(v),\phi(v^{\prime})\in V_{1}$. Moreover, if $N_{C(G)}(a)$ is the open neighborhood of $a\in V(C(G))$, then
    $\phi(\{v,v^{\prime}\})\in N_{C(G)}(\phi(v))\cap N_{C(G)}(\phi(v^{\prime}))$,
    since automorphisms preserve adjacency. Thus, $\{\phi(v),\phi(v^{\prime})\}=\phi(\{v,v^{\prime}\})$ since $G$ is a simple graph. Thus, we have the following:
    $g(\phi\{v,v^{\prime}\})=g(\{v,v^{\prime}\})$ $\implies g(\{\phi(v),\phi(v^{\prime})\})=g(\{v,v^{\prime}\})$
    $\implies f(\{\phi(v),\phi(v^{\prime})\})=f(\{v,v^{\prime}\})$
    $\implies f(\phi\{v,v^{\prime}\})=f(\{v,v^{\prime}\}).$
Thus, $\psi$ preserves $f$. Since all vertices in $V_{1}$ have different colors with respect to $f$, only the identity automorphism of $\overline{G}$ will be the $f$-preserving automorphism of $\overline{G}$. Thus, $\phi(u)=\psi(u)=u$ for all $u\in V_{1}$. By Lemma 2.3, we have $\phi=id_{C(G)}$. ∎
(2). Since $G$ is regular and $G\not\cong K_{n}$, $\overline{G}$ is regular with non-zero degree $\Delta(\overline{G})=n-1-\Delta(G)$. By assumption, $\Delta(\overline{G})\geq\frac{2n}{3}+\frac{23}{6}$. Thus, by Fact 4.3(4), $\overline{G}$ has TCC. Consequently, by Fact 4.3(5), $\chi_{D}^{\prime\prime}(\overline{G})\in\{\Delta(\overline{G})+1,\Delta(\overline{G})+2\}$. Since $\chi_{D}^{\prime\prime}(C(G))\geq\Delta(C(G))+1$, it suffices to show the following.
Claim 4.8.
$\chi_{D}^{\prime\prime}(C(G))\leq\Delta(C(G))+1$.
Proof.
Consider the bipartite subgraph $H=C(G)\backslash E(\overline{G})$ of $G$. We note that $\Delta(G)=\Delta(H)$. There are two cases to be considered.
Case(i): $\chi_{D}^{\prime\prime}(\overline{G})=\Delta(\overline{G})+1$.
Let $f_{1}$ be a proper total coloring of $\overline{G}$ with colors $11,...,\Delta(\overline{G})+$ 1 such that $f_{1}$ is preserved only by $id_{\overline{G}}$. By Fact 4.3(1), there exists a proper edge coloring $f_{2}$ of $H$ with $\Delta(G)$ new colors, say $\Delta(\overline{G})+2,...,\Delta(\overline{G})+\Delta(G)+1$. We recall that $\Delta(\overline{G})+\Delta(G)+1=n$. Define a proper total coloring $f_{3}$ of $C(G)$ with $n$ colors letting $f_{3}(v)=f_{1}(v)$ for all $v\in V(\overline{G})$, $f_{3}(e)=f_{1}(e)$ for all $e\in E(\overline{G})$, $f_{3}(e)=f_{2}(e)$ for all $e\in E(H)$, and if $v^{\prime}=w_{u_{i},u_{j}}\in V_{2}$ is a vertex that subdivides the edge $\{u_{i},u_{j}\}\in E(G)$, then let $f_{3}(v^{\prime})$ be any color from the set $\{1,...,n\}\backslash\{f_{3}(\{w_{u_{i},u_{j}},u_{i}\}),f_{3}(\{w_{u_{i},u_{j}},u_{j}\}),f_{3}(u_{i}),f_{3}(u_{j})\}$, since $n\geq 5$.
Case(ii): $\chi_{D}^{\prime\prime}(\overline{G})=\Delta(\overline{G})+2$.
By the arguments in Case (i), let $f_{1}$ be a proper total coloring of $\overline{G}$ with colors $21,...,\Delta(\overline{G})+$ 2 such that $f_{1}$ is preserved only by $id_{\overline{G}}$ and $f_{2}$ be a proper edge coloring of $H$ with $\Delta(G)$ new colors, say $\Delta(\overline{G})+3,...,\Delta(\overline{G})+2+\Delta(G)$. Now, $\Delta(\overline{G})+2+\Delta(G)=n+1$. Define a proper total coloring $f_{3}$ of $C(G)$ with $n$ colors as follows:
  • $f_{3}(v)=f_{1}(v)$ for all $v\in V(\overline{G})$, and $f_{3}(e)=f_{1}(e)$ for all $e\in E(\overline{G})$,
  • $f_{3}(e)=f_{2}(e)$ for all $e\in E(H)$ such that $f_{2}(e)\neq n+1$,
  • if $e=\{u,v\}$ such that $u\in V_{1}$ and $v\in V_{2}$, and $f_{2}(e)=n+1$, then let $f_{3}(e)$ be any color from the set $\{1,...,\Delta(\overline{G})+2\}\backslash\{f_{3}(u)\}\cup\{f_{3}(e^{\prime}):e^{\prime}\in E(\overline{G}),u\text{ is incident to }e^{\prime}\}$. This is possible since $deg_{\overline{G}}(u)\leq\Delta(\overline{G})$.
  • If $v^{\prime}=w_{u_{i},u_{j}}\in V_{2}$ subdivides the edge $\{u_{i},u_{j}\}\in E(G)$, then we define $f_{3}(v^{\prime})$ as in Case (i).
Since $f_{1}$ is a proper total coloring of $\overline{G}$ that is preserved only by $id_{\overline{G}}$, we have that $f_{3}$ is only preserved by $id_{C(G)}$ following the arguments of Claim 4.7. Consequently, $\chi_{D}^{\prime\prime}(C(G))\leq n=\Delta(C(G))+1$. ∎
This concludes the proof of Theorem 4.5(2). ∎
4.2. Subdivision graphs
Proposition 4.9.
Let $G$ be a connected graph of order $n\geq 5$. Then $\chi^{\prime\prime}_{D}(S(G))\in\{\Delta(S(G))+1,\Delta(S(G))+2\}$. Moreover, if there exists a vertex $x$ of $G$ for which $\phi(x)=x$ for all $\phi\in Aut(G)$, then $\chi^{\prime\prime}_{D}(S(G))$ is $\Delta(S(G))+1$.
Proof.
Since $S(G)$ is a bipartite graph with bipartitions $V_{1}$ and $V_{2}$, $S(G)$ has an edge coloring $g$ with $\Delta(S(G))$ colors according to Fact 4.3 (1). For all $v\in V_{1}$, since $deg_{S(G)}(v)\leq\Delta(S(G))$, we have
$S_{v}=\{1,...,\Delta(S(G))+1\}\backslash\{g(e):v\text{ is incident to }e\}\neq\emptyset$.
Define the total coloring $g_{1}$ of $S(G)$ with $\Delta(S(G))+1$ colors letting $g_{1}(e)=g(e)$ for all $e\in E(S(G))$, $g_{1}(v)$ be any color from $S_{v}$ for all $v\in V_{1}$, and $g_{1}(w)$ be any color out of $\Delta(S(G))+1$ colors different from the color of its neighbors and edges incident on it for all $w\in V_{2}$.
Suppose that there exists a vertex $x$ of $G$ for which $\phi(x)=x$ for all $\phi\in Aut(G)$. In this case, any neighbor of $x$ has to be mapped into itself because every edge incident to $x$ has distinct color. Hence, by induction on the distance from $x$, we can conclude that $g_{1}$ is only preserved by $id_{S(G)}$.
Assume that there are no vertices $x$ of $G$ for which $\phi(x)=x$ for all $\phi\in Aut(G)$. Pick any $x\in V(G)$ and redefine $g_{1}(x)=\Delta(S(G))+2$. By the arguments in Case(i), $g_{1}$ is only preserved by $id_{S(G)}$. ∎
5. AVD-total chromatic number
5.1. Central graphs of regular graphs
Theorem 5.1.
If $G$ is a regular connected graph with order $n\geq 5$ then the following holds:
  1. (1)
    If $n$ is even then $C(G)$ is an AVD Type 2 graph.
  2. (2)
    If $n$ is odd, then $C(G)$ satisfies AVD-TCC.
Proof.
We note that $\Delta(C(G))=n-1$.
[…]
($\ast$) the color class $C_{C(G)}(u_{i})$ contains all colors of $\{m_{i,j}:1\leq j\leq\Delta(C(G))+1\}$ and $m_{i,\Delta(C(G))+2}$ is different for each $i$, $1\leq i\leq n$.
Since $M(k)$ is anti-circulant (see Def. 4.1) and ()($\ast$)( ∗ ) hold, we have $C_{C(G)}(u_{i})\neq C_{C(G)}(u_{j})$ for any two distinct vertices $u_{i}$ and $u_{j}$ in $V_{1}$. Since for all $u_{i}\in V_{1}$, $deg_{C(G)}(u_{i})\neq 2$ and for all $v_{j}\in V_{2}$, $deg_{C(G)}(v_{j})=2$, we have $C_{C(G)}(u_{i})\neq C_{C(G)}(v_{j})$ for each $u_{i}\in V_{1}$ and $v_{j}\in V_{2}$. Thus, $\pi$ is an AVD-total coloring of $C(G)$. Since the vertices of maximum degree are adjacent in $C(G)$, we have $\chi_{a}^{\prime\prime}(C(G))=\Delta(C(G))+2$.
[…]
5.2. Subdivision graphs
Recall that $S(G)$ is the subdivision graph of $G$, $V_{1}=V(G)$, and $V_{2}$ is the set of all subdivided vertices. We note that every bipartite graph $G$ has ($\Delta(G)+2$)-AVD total coloring where $\Delta(G)$ colors are used to color the edges of $G$ and two new colors are used to color the vertices of the bipartitions.
Theorem 5.2.
([7, Theorem 2.3]) If $G$ is a bipartite graph, then $\chi^{\prime\prime}_{a}(G)\leq\Delta(G)+2$.
Since $S(G)$ is a bipartite graph, $\chi^{\prime\prime}_{a}(S(G))\leq\Delta(S(G))+2$ by the above result. However, we observe that if $\Delta(G)\geq 5$, then $\chi^{\prime\prime}_{a}(S(G))=\Delta(S(G))+1$.
Theorem 5.3.
If $G$ is a connected graph and $\Delta(G)\geq 5$ then $S(G)$ is $AVD$-Type 1.
Proof.
Since $\Delta(G)\geq 5$, we have $\Delta(S(G))=\Delta(G)$. Since $\Delta(G)+1=\Delta(S(G))+1\leq\chi^{\prime\prime}_{a}(S(G))$, it suffices to show that $\chi^{\prime\prime}_{a}(S(G))\leq\Delta(G)+1$. We construct an AVD-total-coloring $f$ of $S(G)$ with colors from the set $S=\{1,...,\Delta(G)+1\}$. By Fact 4.3 (1), there is a proper edge coloring $g$ of $S(G)$ with colors from the set $\{1,...,\Delta(G)\}$ as $S(G)$ is a bipartite graph. Let $f(e)=g(e)$ for any $e\in E(S(G))$. Let $f(v)=\Delta(G)+1$ for all $v\in V_{1}$. Next, we color all the vertices of $V_{2}$ so that if $v^{\prime}=w_{v,w}\in V_{2}$ then $C_{S(G)}(v^{\prime})\neq C_{S(G)}(v)$ and $C_{S(G)}(v^{\prime})\neq C_{S(G)}(w)$ where $C_{S(G)}(x)$ is the color class of $x$ in $S(G)$. Fix $v^{\prime}=w_{v,w}\in V_{2}$.
  1. Case A: $deg_{S(G)}(v)\neq 2$ and $deg_{S(G)}(w)\neq 2$. We note that $f(w)=f(v)=\Delta(G)+1$. Let
    $R_{v^{\prime}}=\{1,...,\Delta(G)+1\}\backslash\{f(v),f(\{v,v^{\prime}\}),f(\{w,v^{\prime}\})\}$.
  2. Case B: $deg_{S(G)}(v)=deg_{S(G)}(w)=2$. Let $N_{S(G)}(v)=\{w_{v,w},w_{v,w_{1}}\}$, $N_{S(G)}(w)=\{w_{v,w},w_{v,w_{2}}\}$, and
    $R_{v^{\prime}}=\{1,...,\Delta(G)+1\}\backslash\{f(v),f(\{v,v^{\prime}\}),f(\{w,v^{\prime}\}),f(\{v,w_{v,w_{1}}\}),f(\{w,w_{v,w_{2}}\})\}$.
  3. Case C: $deg_{S(G)}(v)=2$ and $deg_{S(G)}(w)\neq 2$. Assume $N_{S(G)}(v)$ as in Case B. Let
    $R_{v^{\prime}}=\{1,...,\Delta(G)+1\}\backslash\{f(v),f(\{v,v^{\prime}\}),f(\{w,v^{\prime}\}),f(\{v,w_{v,w_{1}}\})\}$.
In each of the above cases, $R_{v^{\prime}}\neq\emptyset$ since $\Delta(G)\geq 5$. Let $f(v^{\prime})$ be any color from $R_{v^{\prime}}$.
Finally, $f$ is a total coloring of $S(G)$ and since $C_{S(G)}(v)\neq C_{S(G)}(w_{v,w})$ for any fixed edge $\{v,w_{v,w}\}\in E(S(G))$, $f$ is an AVD-total coloring of $S(G)$. ∎
5.3. Central graphs of join of two graphs
Fact 5.4.
The following holds:
  1. (1)
    (Zhang et al. [37]) Let $K_{n}$ be a complete graph of order $n$, $n\geq 3$, then
    $\chi^{\prime\prime}_{a}(K_{n})=\begin{cases}n+1&\text{if }\,n\equiv 0\text{(mod 2)},\\ n+2&\text{if }\,n\equiv 1\text{(mod 2)}.\end{cases}$
  2. (2)
    (Panda et al. [28]) If $G\in\{P_{n},C_{n}\}$, then $\chi^{\prime\prime}_{a}(C(G))=n+1$ if $n$ is even and $\chi^{\prime\prime}_{a}(C(G))\leq n+2$ otherwise.
  3. (3)
    (Panda et al. [28]) $\chi^{\prime\prime}_{a}(C(K_{1,n}))=n+2$ and $\chi^{\prime\prime}_{a}(C(K_{n}))=n$ for any integer $n\geq 1$.
Proposition 5.5.
The following holds:
  1. (1)
    If $C(G_{1})$ and $C(G_{2})$ satisfy AVD-TCC, then $C(G_{1}+G_{2})$ satisfies AVD-TCC.
  2. (2)
    The graph $C(K_{m,n})$ satisfies AVD-TCC for any $m,n\geq 2$.
  3. (3)
    If $G_{1},G_{2}\in\{P_{n},C_{n},K_{n}\}$, then $C(G_{1}+G_{2})$ satisfies AVD-TCC for any integer $n\geq 2$.
  4. (4)
    If $G_{1}$ and $G_{2}$ have the same order, then $C(G_{1}+G_{2})$ satisfies AVD-TCC.
Proof.
The proofs are straightforward using some coloring arguments and applying the results mentioned in Fact 5.4, and thus we provided the details in the Appendix. ∎
6. Total dominator chromatic number
Theorem 6.1.
Let $G$ be a connected graph of order $n\geq 5$. The following holds:
  1. (1)
    If $\Delta(G)\leq n-3$, then $\chi^{t}_{d}(\overline{G})\leq\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
  2. (2)
    If $G$ is a tree, then $\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
Moreover, the upper bounds of $\chi^{t}_{d}(C(G))$ mentioned in (1) and (2) are sharp.
Proof.
(1). Let $V_{1}=V(G)=\{v_{1},...,v_{n}\}$. Let $V_{2}=V(C(G))\backslash V(G)=\{w_{i,j}:\{v_{i},v_{j}\}\in E(G)\}$ be the set of subdivided vertices and $f=(X_{1},...,X_{l})$ be a TDC (i.e., total dominator coloring) of $C(G)$. Then, for every $v_{i}\in V_{1}$, there exists $k\leq l$ such that $X_{k}\subseteq N_{C(G)}(v_{i})$. Define $g=(W_{1},...,W_{l})$ where $W_{i}=X_{i}\backslash V_{2}$ for all $1\leq i\leq l$. Clearly, $g$ is a proper vertex coloring of $\overline{G}$.
Case (i): If for every $v_{i}\in V_{1}$, there exists $k\leq l$ such that $X_{k}\subseteq N_{C(G)}(v_{i})$ and $W_{k}=X_{k}\cap V_{1}\neq\emptyset$, then $W_{k}\subseteq N_{\overline{G}}(v_{i})$, and so $g$ is a TDC of $\overline{G}$.
[…]
Figure 4. In the figure above, shaded regions contain vertices belonging to $V_{2}$.
Case (ii): Assume that there exists $v_{i}\in V_{1}$ such that $\bigcup_{1\leq t\leq r}X_{k_{t}}\subseteq N_{C(G)}(v_{i})\cap V_{2}$ for some $k_{1},...,k_{r}\leq l$. Since $\Delta(G)<n-2$, this implies $N_{C(G)}(v_{i})\cap V_{1}\neq\emptyset$. Let $N_{C(G)}(v_{i})\cap V_{1}=\{m_{1},...,m_{p}\}$ for some positive integer $p$. For every $1\leq j\leq p$, assume that $m_{j}\in X_{t_{j}}$ for some $t_{j}\leq l$. If for some $1\leq j\leq p$, we have $X_{t_{j}}\cap V_{1}\subseteq N_{C(G)}(v_{i})$, then we are done. Otherwise, assume that for all $1\leq j\leq p$, we have $X_{t_{j}}\cap V_{1}\not\subseteq N_{C(G)}(v_{i})$. Since $m_{j}\in X_{t_{j}}\cap V_{1}\cap N_{C(G)}(v_{i})$, we have $|X_{t_{j}}\cap V_{1}|\geq 2$. Specifically, we have $X_{t_{1}}\cap V_{1}\not\subseteq N_{C(G)}(v_{i})$ and $|X_{t_{1}}\cap V_{1}|\geq 2$. Note that $X_{k_{1}}\subseteq V_{2}$, and so $W_{k_{1}}=\emptyset$. Redefine $W_{k_{1}}=\{m_{1}\}$ and $W_{t_{1}}=X_{t_{1}}\backslash(V_{2}\cup\{m_{1}\})$. Then, $W_{k_{1}}\subseteq N_{\overline{G}}(v_{i})$. If we do the above procedure on all $v_{i}\in V_{1}$ such that $\bigcup_{1\leq t\leq r}X_{k_{t}}\subseteq N_{C(G)}(v_{i})\cap V_{2}$ for some $k_{1},...,k_{r}\leq l$, then $g=(W_{1},...,W_{l})$ is a TDC of $\overline{G}$.
Claim 6.2.
$\chi^{t}_{d}(C(G))\leq n$. Moreover, the bound $\chi^{t}_{d}(C(G))\leq n$ is sharp.
Proof.
Let $T_{n}=\{v_{i}:\{v_{i},v_{n}\}\in E(G)\}$. Since $G$ is connected, $T_{n}\not=\emptyset$. Pick $k\in\{i:v_{i}\in T_{n}\}$. Let $X_{i}=\{v_{i}\}$ for all $i\in\{1,...,n-1\}\backslash\{k\}$, $X_{n}=V_{2}$, and $X_{k}=\{v_{k},v_{n}\}$.111i.e., we color all the vertices of $V_{2}$ with $n$, assign the colors $11,...,n-$ 1 to $v_{1},...,v_{n-1}$, and color $v_{n}$ with $k$. For the following reasons, $g=(X_{1},...,X_{n})$ is a TDC of $G$:
  1. (1)
    $X_{k}\subseteq N_{C(G)}(w_{v_{n},v_{k}})$, and for any $v\in V_{2}\backslash\{w_{v_{n},v_{k}}\}$, there exists $t\in\{1,...,n-1\}$ where $X_{t}\subseteq N_{C(G)}(v)$.
  2. (2)
    Fix $x\in V_{1}\backslash\{v_{n},v_{k}\}$.
    1. (a)
      If $\{v_{n},v_{k}\}\subseteq N_{C(G)}(x)$, then $X_{k}\subseteq N_{C(G)}(x)$.
    2. (b)
      If exactly one of $v_{n}$ and $v_{k}$ is adjacent to $x$ in $C(G)$, then $x$ is adjacent to some vertex $v_{t}$ in $V_{1}\backslash\{v_{n},v_{k}\}$ since $\Delta(G)\leq n-3$. Thus, $X_{t}\subseteq N_{C(G)}(x)$.
    3. (c)
      If neither $v_{n}$ nor $v_{k}$ is adjacent to $x$ in $C(G)$, then we follow step (b).
  3. (3)
    Let $y\in\{v_{n},v_{k}\}$. Following step (b), there exists $t\in\{1,...,n-1\}$ such that $X_{t}\subseteq N_{C(G)}(y)$.
We show that the bound $\chi^{t}_{d}(C(G))\leq n$ is sharp. Define
$G=K_{n}\backslash(\{\{v_{i},v_{i+1}\}\in E(K_{n}):1\leq i\leq n-1\}\cup\{\{v_{n},v_{1}\}\})$
where the vertices $v_{1},...,v_{n}$ of $K_{n}$ are arranged in a clockwise order (see Fig. 5). Then $\Delta(G)=n-3$. It is enough to show that $\chi^{t}_{d}(C(G))\geq n$.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$
Figure 5. Graph $G$ when $n=6$.
We note that for any TDC $f=(X_{1},...,X_{l})$ of $C(G)$, |$|\{X_{i}:X_{i}\subseteq V_{1},|X_{i}|=1\}|\geq n-2$. Otherwise, the following cases can occur.
Case (i): There exist two color classes of two elements, say $X_{1}=\{v_{i},v_{j}\}$ and $X_{2}=\{v_{k},v_{t}\}$. Observe that both $v_{i}\in N_{C(G)}(v_{k})\cap N_{C(G)}(v_{t})$ and $v_{j}\in N_{C(G)}(v_{k})\cap N_{C(G)}(v_{t})$ can’t happen together by the definition of $G$, since $n\geq 5$. Without loss of generality, assume that $v_{i}\not\in N_{C(G)}(v_{k})$. Thus, $v_{i}$ is adjacent to $v_{k}$ in $G$. Then $N_{C(G)}(v_{i})\cap N_{C(G)}(v_{k})=\{w_{i,k}\}$ and $X_{m}\not\subseteq N_{C(G)}(w_{i,k})$ for each $1\leq m\leq l$; a contradiction to the fact that $f$ is a TDC of $C(G)$.
Case (ii): There exists a color class of $m$ elements, say $X_{1}=\{v_{i_{1}},...,v_{i_{m}}\}$, where $m\geq 3$. We observe that for any $k,l\in\{i_{1},...i_{m}\}$ where $k\neq l$, we have $\{v_{k},v_{l}\}\in E(G)$. Then $N_{C(G)}(v_{l})\cap N_{C(G)}(v_{k})=\{w_{l,k}\}$, $X_{1}\not\subseteq N_{C(G)}(w_{l,k})$ and the contradiction follows from the arguments of Case(i).
Therefore, |$|\{f(v_{i}):1\leq i\leq n\}|\geq n-1$. Since at least one new color is needed to color some vertices in $C=\{w_{i,j}:\{v_{i},v_{j}\}\in E(G)\}$, we have $l\geq n$. ∎
(2). In view of (1), it suffices to show that if $\Delta(T)\geq n-3$, then $\chi^{t}_{d}(C(T))\leq n$. The proof is trivial, and hence we sketched the details of the proof in the Appendix. ∎
7. Open Problems
The following questions remain open.
Question 7.1.
Does AVD-TCC hold for the central graph of any graph?
Question 7.2.
Classify all graphs $G$ (other than regular graphs) so that $C(G)$ is a Type 1 graph.
Funding The research of the first author was supported by the EKÖP-24-4-II-ELTE-996 University Excellence scholarship program of the Ministry of Culture and Innovation from the source of the National Research, Development and Innovation fund.

### v2

@@
-On distinguishing coloring and some variants of proper coloring of graphs derived from subdivision operations
+Automorphism groups and Distinguishing Colorings of Central and Middle Graphs
 Abstract.
-Let $G$ be a simple, finite, connected, and undirected graph, and $T$ be a finite tree.
+Let $G$ be a simple, finite, connected, and undirected graph.
 The middle graph $M(G)$ of $G$ is obtained from the subdivision graph $S(G)$ after joining pairs of subdivided vertices that lie on adjacent edges of $G$ and the central graph $C(G)$ of $G$ is obtained from $S(G)$ after joining all non-adjacent vertices of $G$.
-We show that if the order of $G$ is at least $4$, then $Aut(G)\cong Aut(C(G))\cong Aut(M(G))$(as abstract groups) and apply these results to obtain new sharp upper bounds of the distinguishing number and the distinguishing index of $C(G)$ and $M(G)$ inspired by an algorithm due to Kalinowski, Pilśniak, and Woźniak from 2016.
-Furthermore, we study the total distinguishing chromatic number of $C(G)$ and $S(G)$, use Latin squares to verify the AVD-total coloring conjecture for central graphs of regular graphs and some other classes of graphs (which is a partial progress towards answering an open question of Panda, Verma, and Keerti), and obtain new bounds of the total dominator chromatic number of $C(G)$ and $C(T)$.
+We show that if the order of $G$ is at least $4$, then $Aut(G)\cong Aut(C(G))\cong Aut(M(G))$(as abstract groups) and apply these results to obtain new upper bounds of the distinguishing number and the distinguishing index of $C(G)$ and $M(G)$ inspired by an algorithm by Kalinowski, Pilśniak, and Woźniak from 2016.
 Key words and phrases:
-Automorphism groups, distinguishing coloring, total dominator coloring, adjacent vertex distinguishing total coloring, central graph, subdivision graph
+Automorphism group, distinguishing coloring, central graph, middle graph, subdivision graph
 2020 Mathematics Subject Classification:
@@
 Corresponding author.
+In this revised version, all results except Propositions 3.2 and 3.3 were obtained and written by the first author.
+The ideas underlying Propositions 3.2 and 3.3 originate from an earlier version of the manuscript by all authors, where they appeared in the form of diagrams.
+The propositions are formalized and proved here by the first author.
 1. Introduction
-Let $Aut(G)$ be the automorphism group of a graph $G=(V(G),E(G))$, $id_{G}$ be the trivial automorphism of $G$, and $f$ be a vertex coloring or an edge coloring or a total coloring (that is, a coloring of its vertices and edges) of $G$.
-An automorphism $\phi$ of $G$ preserves $f$ if the following holds: (1) $f(x)=f(\phi(x))$ for all $x\in V(G)$ if $f$ is a vertex coloring, (2) $f(e)=f(\phi(e))$ for all $e\in E(G)$ if $f$ is an edge coloring, (3) $f(x)=f(\phi(x))$ for all $x\in V(G)\cup E(G)$ if $f$ is a total coloring.
-1.1. Distinguishing number and distinguishing index
-Albertson and Collins [1] introduced the distinguishing number $D(G)$ of a graph $G$ as the least integer $d$ such that $G$ admits a vertex coloring with $d$ colors that is only preserved by $id_{G}$ and Kalinowski and Pilśniak [16] introduced the distinguishing index $D^{\prime}(G)$ of $G$ as the least integer $d$ such that $G$ admits an edge coloring with $d$ colors that is only preserved by $id_{G}$.
-In [4, Theorems 3.8 and 4.3], Alikhani and Soltani proved that if $G$ is a connected graph of order $n\geq 3$, then $D(S(G))\leq\lceil\sqrt{\Delta(G)}\rceil$ and $D^{\prime}(S(G))\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$ where $\Delta(G)$ is the maximum degree of $G$.
-Hamada and Yoshimura [13] and Vernold [31] introduced the concept of middle graphs and central graphs, respectively; classes of graphs derived from subdivision graphs.
-Since then, the notion of central graphs and middle graphs has been intensively studied in graph theory; see, for example, [5, 9, 11, 12, 18, 19, 23, 28].
-Results: Let $G$ be a connected graph of order $n$. The following holds:
-  1. (1)
-    If $n\geq 3$, then $D(M(G))\leq\Delta(G)$,
-  2. (2)
-    $D^{\prime}(M(G))\leq 3$ for any positive integer $n$, and
-  3. (3)
-    If $n\geq 4$, then $D^{\prime}(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$, and $D(C(G))\leq\lceil\sqrt{\Delta(G)}\rceil$.
-Moreover, the bounds mentioned in (1) and (3) are sharp (see Theorems 3.2, 3.4, 3.6).
-1.2. Total distinguishing chromatic number
-A proper total coloring of a graph $G$ is a total coloring of $G$ so that no adjacent vertices, no adjacent edges, and no incident vertices and edges get the same color.
-The least number of colors needed for a proper total coloring of $G$ is called its total chromatic number, denoted by $\chi^{\prime\prime}(G)$.
-Behzad [6] and Vizing [33] independently posed the following conjecture.
-Conjecture 1.1.
-(Total-Coloring Conjecture (TCC)). For any graph $G$, $\chi^{\prime\prime}(G)\leq\Delta(G)+2$.
-Recently, Panda et al. [28, Theorem 3.1] proved that TCC holds for the central graph of any graph of order $n\geq 5$. We observe an alternative short proof of the result using the following lemma.
-Lemma 1.2.
-([36, Lemma 2.4]) If $G$ is a graph of order $n$ that contains an independent set $S$ such that $|S|\leq n-\Delta(G)-1$, then $G$ satisfies TCC.
-Theorem 1.3.
-(Panda–Verma–Keerti;[28]) TCC holds for the central graph of any graph.
-Short proof of Theorem 1.3.
-Let $C(G)$ be the central graph of a graph $G=(V(G),E(G))$ where $V_{1}=V(G)$, $V_{2}=V(C(G))\backslash V(G)$, $|V_{1}|=n$, and $|V_{2}|=r$.
-If $n\geq 3$, then the order of $C(G)$ is $n+r$, $\Delta(C(G))=n-1$, and $C(G)$ contains an independent set $V_{2}$ of size $r=(n+r)-\Delta(C(G))-1$.
-By Lemma 1.2, $C(G)$ satisfies TCC.
-If $n\leq 2$, then it is trivial to verify.
-∎
-We say $G$ is of Type $j$ if $\chi^{\prime\prime}(G)=\Delta(G)+j$ for $j\in\{1,2\}$.
-In [15], Kalinowski, Pilśniak, and Woźniak introduced the total distinguishing chromatic number $\chi^{\prime\prime}_{D}(G)$ of a graph $G$ as the least integer $d$ such that $G$ has a proper total coloring with $d$ colors that is preserved only by $id_{G}$.
-We study the total distinguishing chromatic number of central and subdivision graphs.
-Results: Let $G$ be a connected graph of order $n$. The following holds:
-  1. (1)
-    (Proposition 4.4) $\Delta(C(G))+1\leq\chi^{\prime\prime}(C(G))\leq\chi^{\prime\prime}_{D}(C(G))\leq\Delta(C(G))+2$.
-  2. (2)
-    (Theorem 4.5(1)) If $G$ is regular and $n\geq 5$ is odd, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
-  3. (3)
-    (Theorem 4.5(2)) If $G$ is regular and a non complete graph, $n\geq 5$ is even, and $\Delta(G)\leq\frac{2n-29}{6}$, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
-  4. (4)
-    (Proposition 4.9) $\chi^{\prime\prime}_{D}(S(G))$ is either $\Delta(S(G))+1$ or $\Delta(S(G))+2$.
-1.3. AVD-total chromatic number
-An adjacent vertex distinguishing (AVD) total coloring of a graph $G$ with edge set $E(G)$ is a proper total coloring $f$ such that for any edge $\{u,v\}\in E(G)$, we have $C_{G}(u)\neq C_{G}(v)$, where $C_{G}(u)=\{f(u)\}\cup\{f(\{u,v\}):\{u,v\}\in E(G)\}$ is the color class of $u$.
-The adjacent vertex distinguishing (AVD)-total chromatic number $\chi^{\prime\prime}_{a}(G)$ of $G$ is the minimum integer $k$ such that there exists an AVD-total coloring of $G$ using $k$ colors.
-Zhang et al.
-[37] introduced AVD-total-colorings and posed the following conjecture.
-Conjecture 1.4.
-(AVD-TCC). For any graph $G$, $\chi^{\prime\prime}_{a}(G)\leq\Delta(G)+3$.
-We say $G$ is of AVD-Type $i$ if $\chi^{\prime\prime}_{a}(G)=\Delta(G)+i$ for $i\in\{1,2,3\}$.
-The Conjecture 1.4 for general graphs $G$ with $\Delta(G)=3$ has been verified by Wang [34] and Chen [7].
-Lu et al.
-[22] validated the conjecture for all graphs $G$ with $\Delta(G)=4$.
-The Conjecture 1.4 is also known to be true for many families of graphs such as 4-regular graphs [26], hypercubes [8], complete equipartite graphs [24], indifference graphs [27], split graphs [32], and several other classes of graphs.
-Panda et al.
-[28] obtained the AVD-total chromatic number of central graphs for some known families of graphs and mentioned the following open question (see [28, Section 5]).
-Question 1.5.
-(Panda et al. [28]) Does AVD-TCC hold for the central graph of any graph?
-In [28, Section 5], it was mentioned that it would be interesting to solve the AVD-total coloring classification problem on the central graph of any graph.
-We have made partial progress in this direction.
-Results: The following holds:
-  1. (1)
-    (Theorem 5.1) If $G$ is a regular graph of even order $n\geq 5$, then $C(G)$ is AVD-Type 2 and if $G$ is a regular graph of order $n\geq 5$, then $C(G)$ satisfies AVD-TCC.
-  2. (2)
-    (Theorem 5.3) If $\Delta(G)\geq 5$, then the subdivision graph $S(G)$ of $G$ is AVD-Type 1.
-  3. (3)
-    (Proposition 5.5 (4)) If two graphs $G_{1}$ and $G_{2}$ have the same order and $G_{1}+G_{2}$ is the join of $G_{1}$ and $G_{2}$, then $C(G_{1}+G_{2})$ satisfies AVD-TCC.
-1.4. Total dominator chromatic number
-The notion of total domination number was introduced by Cockayne, Dawes and Hedetniemi in [10] and has been studied for central graphs and middle graphs, see [9, 18, 19, 20, 21].
-In particular, Kazemnejad and Moradi [19] and Chen et al.
-[9] studied sharp upper bounds for the total domination number of central graphs of connected graphs and trees.
-Motivated by the relation between the chromatic number and the total domination number, the notion of the total dominator chromatic number was studied in [17].
-A total dominator coloring, briefly TDC, of a graph $G$ is a proper vertex coloring of $G$ in which each vertex of $G$ is adjacent to every vertex of some color class.
-The total dominator chromatic number $\chi^{t}_{d}(G)$ of $G$ is the minimum number of color classes in a TDC of $G$.
-The total dominator chromatic number was studied for middle graphs by Kazemnejad et al.
-[20], $k$-subdivision of graphs by Alikhani et al.
-[2], and central graphs by Kazemnejad et al.
-[18].
-In [18, Theorem 2.1], Kazemnejad and Kazemi proved that the following bounds are sharp if $G$ is a connected graph of order $n\geq 2$:
-  • $\lfloor 2n/3\rfloor+1\leq\chi^{t}_{d}(C(G))\leq\Delta(C(G))+2$ if $\Delta(G)\leq n-2$.
-  • $\chi^{t}_{d}(C(G))\leq n+\lceil n/2\rceil$ if $G$ has a Hamiltonian path.
-Results: Let $G$ be a connected graph of order $n\geq 5$. The following holds:
-  1. (1)
-    If $\Delta(G)\leq n-3$ and $\overline{G}$ is the complement of $G$, then $\chi^{t}_{d}(\overline{G})\leq\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
-  2. (2)
-    If $G$ is a tree, then $\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
-Moreover, the upper bounds of $\chi^{t}_{d}(C(G))$ mentioned in (1) and (2) are sharp (see Theorem 6.1).
-1.5. Tables of results
+All graphs considered in this paper are finite, simple, connected, and undirected.
+Let $G=(V(G),E(G))$ be a connected graph with the vertex set $V(G)$ and the edge set $E(G)$.
+Let $\Delta(G)$ represent the maximum degree of $G$.
 Let $C_{n}$ be a cycle graph of $n$ vertices, $P_{n}$ be a path of $n$ vertices, $K_{n}$ be a complete graph of $n$ vertices, and $K_{n,m}$ be a complete bipartite graph with bipartitions of sizes $n$ and $m$.
 A claw is the graph $K_{1,3}$.
-The degree of a vertex $v\in V(G)$ in $G=(V(G),E(G))$, denoted by $deg_{G}(v)$, is the number of edges that emerge from $v$.
+The degree of a vertex $v$ in $G$, denoted by $deg_{G}(v)$, is the number of edges that emerge from $v$.
 We say $G$ is a regular graph if all vertices of $G$ have the same degree.
 The degree of a regular graph $G$, denoted by $d(G)$, is the same as the maximum degree $\Delta(G)$ and the minimum degree $\delta(G)$.
-Let $N_{G}(v)=\{u\in V(G):\{u,v\}\in E(G)\}$ for any $v\in V(G)$.
-Graph $D^{\prime}(H)$ $D(H)$
-$H=C(G),n\geq 4$ $\leq\lceil\sqrt{\Delta(G)}\rceil$ (Thm. 3.2) $\leq\lceil\sqrt{\Delta(G)}\rceil$ (Thm. 3.4)
-$H=M(G),n\geq 3$ $\leq 3$ (Thm. 3.6(2)) $\leq\Delta(G)$ (Thm. 3.6(1))
-$H=S(G)$ $\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$ ([4, Thm. 4.3]) $\leq\lceil\sqrt{\Delta(G)}\rceil$ ([4, Thm. 3.8])
-Graph $\chi^{\prime\prime}_{a}(H)$ $\chi^{\prime\prime}_{D}(H)$
-$H=S(G),\Delta(G)\geq 5$ $\Delta(H)+1$ (Thm. 5.3) $\leq\Delta(H)+2$ (Prop. 4.9)
-$H=C(G),n\equiv 0$(mod 2), $G\not\cong K_{n}\text{ is regular}$, $n\geq 5$, and $\Delta(G)\leq\frac{2n-29}{6}$ $\Delta(H)+2$ (Thm. 5.1) $\Delta(H)+1$ (Thm. 4.5)
-[…]
-$H=C(K_{m,r})$ $\leq\Delta(H)+3$ (Prop. 5.5(2)) $\leq\Delta(H)+2$ (Prop. 4.4)
-$H=C(G_{1}+G_{2})$, $G_{1},G_{2}\in\{P_{n},C_{n},K_{n},K_{n,r}\}$ $\leq\Delta(H)+3$ (Prop. 5.5(3)) $\leq\Delta(H)+2$ (Prop. 4.4)
-$H=C(H_{1}+H_{2})$, order of $H_{1}$ = order of $H_{2}$ $\leq\Delta(H)+3$ (Prop. 5.5(4)) $\leq\Delta(H)+2$ (Prop. 4.4)
-Graph $\chi^{t}_{d}(H)$
-$H=C(G),n\geq 5,\Delta(G)\leq n-3$ $\geq\chi^{t}_{d}(\overline{G})$, $\leq\Delta(H)+1$ (Thm. 6.1(1))
-$H=C(T),T\text{ is a tree, }\text{order of }T\geq 5$ $\leq\Delta(H)+1$ (Thm. 6.1(2))
-$H=M(G),n\geq 2,G\text{ has $m$ edges}$ $\leq n+m-1$ ([20, Thm. 2.3])
-$H=S(G),G\text{ has $m$ edges}$ $\leq 2m$ ([2, Thm. 2.2])
-$H=C(G),n\geq 3,\Delta(G)\leq n-2$ $\leq\Delta(H)+2$ ([18, Thm. 2.1])
-$H=C(G),n\geq 2,\text{ $G$ has a Hamiltonian path}$ $\leq n+\lceil n/2\rceil$ ([18, Thm. 2.1])
-In the above tables, we list the results obtained in this manuscript, where $n$ denotes the order of $G$.
+Let $N_{G}(v)=\{u\in V(G):uv\in E(G)\}$ for any $v\in V(G)$.
+The middle graph $M(G)$ and the central graph $C(G)$, introduced by Hamada and Yoshimura [8] and Vernold [18], respectively, are classes of graphs derived from subdivision graphs.
+These graph constructions have been extensively studied; see, for example, [4, 5, 6, 7, 11, 12, 13, 15].
+Definition 1.1.
+The subdivision graph $S(G)$ of $G$ is obtained from $G$ by putting a new vertex in the middle of every edge of $G$.
+For each edge $xy\in E(G)$, introduce a new vertex $w_{xy}$ (called the subdivision vertex corresponding to the edge $xy$) and replace the edge $xy$ by the two edges $xw_{xy}$ and $w_{xy}y$.
+Thus, $V(S(G))=V(G)\cup\{w_{xy}:xy\in E(G)\},$ and $E(S(G))=\{xw_{xy},\,w_{xy}y:xy\in E(G)\}.$ The middle graph $M(G)$ of $G$ is the graph whose vertex set is $V(G)\cup E(G)$, and two vertices $x,y$ in the vertex set of $M(G)$ are adjacent in $M(G)$ in the case where one of the following holds:
+  1. (1)
+    $x,y\in E(G)$ and $x,y$ are adjacent in $G$;
+  2. (2)
+    $x\in V(G),y\in E(G)$, and $x,y$ are incident in $G$.
+The central graph $C(G)$ of $G$ with order $n$ and size $m$ is a graph of order $n+m$ and size $n\choose{2}$ + $m$ that is obtained from $S(G)$ by adding the edges in $E(\overline{G})$, where $\overline{G}$ is the complement graph of $G$.
+Definition 1.2.
+Let $Aut(G)$ be the automorphism group of $G$, $id_{G}$ be the trivial automorphism of $G$, and $f$ be a vertex coloring or an edge coloring or a total coloring (that is, a coloring of its vertices and edges) of $G$.
+An automorphism $\phi$ of $G$ preserves $f$ if the following holds:
+  1. (1)
+    $f(x)=f(\phi(x))$ for all $x\in V(G)$ if $f$ is a vertex coloring,
+  2. (2)
+    $f(e)=f(\phi(e))$ for all $e\in E(G)$ if $f$ is an edge coloring,
+The notions of distinguishing number and distinguishing index were introduced by Albertson and Collins [1] and Kalinowski and Pilsniak [10], respectively.
+Definition 1.3.
+The distinguishing number (index) of a graph $G$, denoted by $D(G)$ ($D^{\prime}(G)$), is the least integer $d$ such that $G$ admits a vertex coloring (edge coloring) with $d$ colors that is preserved only by $id_{G}$.
+1.1. Motivation
+Alikhani and Soltani [3, Theorems 3.8 and 4.3] studied the distinguishing number and distinguishing index of subdivision graphs and proved that if $G$ is a connected graph of order $n\geq 3$, then
+$D(S(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil\quad\text{and}\quad D^{\prime}(S(G))\leq\min\{r:r^{2}-r\geq D^{\prime}(G)\},$
+where $\Delta(G)$ denotes the maximum degree of $G$.
+Kalinowski–Pilśniak–Woźniak [9] investigated the total distinguishing chromatic number of graphs.
+Inspired by the results of Alikhani and Soltani [3] and the results of Kalinowski–Pilśniak–Woźniak [9], we investigate the distinguishing parameters of middle and central graphs.
+1.2. Main Results
+Let $G$ be a connected graph of order $n$. We prove the following:
+  1. (1)
+    (Theorem 2.9) If $n\geq 4$, then $Aut(G)\cong Aut(C(G))\cong Aut(M(G))$.
+  2. (2)
+    (Theorem 3.10(1)) If $n\geq 3$, then $D(M(G))\leq\Delta(G)$;
+  3. (3)
+    (Theorem 3.10(2)) $D^{\prime}(M(G))\leq 3$;
+  4. (4)
+    (Theorem 3.4) If $n\geq 4$, then $D^{\prime}(C(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil.$
+  5. (5)
+    (Theorem 3.8) If $n\geq 4$, then $D(C(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil.$
 2. Automorphism groups
-Definition 2.1.
-The subdivision graph $S(G)$ of a graph $G=(V(G),E(G))$ is a bipartite graph with the set of vertices $V(S(G))=V_{1}\cup V_{2}$ where $V_{2}=\{w_{v,v^{\prime}}:w_{v,v^{\prime}}\text{ subdivide the edge }\{v,v^{\prime}\}\in E(G)\}$ and $V_{1}=V(G)$, and the set of edges $E(S(G))=\{\{v,w_{v,v^{\prime}}\},\{w_{v,v^{\prime}},v^{\prime}\}\}:\{v,v^{\prime}\}\in E(G)\}$.
-Definition 2.2.
-The central graph $C(G)$ of a graph $G=(V(G),E(G))$ of order $n$ and size $m$ is a graph of order $n+m$ and size ($n\choose{2}$ + $m$ that is obtained by subdividing each edge of $G$ exactly once and joining all the non-adjacent vertices of $G$ in $C(G)$.
-In particular, $C(G)$ is a graph with the set of vertices $V(C(G))=V(S(G))$ and the set of edges $E(C(G))=E(\overline{G})\cup\{\{v,w_{v,v^{\prime}}\},\{w_{v,v^{\prime}},v^{\prime}\}\}:\{v,v^{\prime}\}\in E(G)\}$ where $\overline{G}$ is the complement of $G$.
-In this section, we sketch some lemmas concerning $Aut(C(G))$ and $Aut(M(G))$ where $M(G)$ is the middle graph of $G$. Applying these lemmas, we will prove Theorems 3.2, 3.4, and 4.5.
+In this section, we sketch some lemmas concerning $Aut(C(G))$ and $Aut(M(G))$ where $M(G)$ is the middle graph of $G$.
+Applying these lemmas, we will prove Theorems 3.4 and 3.8 and propositions 3.2, 3.3, and 3.6.
+Recall that for each edge $xy\in E(G)$, $w_{xy}$ denotes the subdivision vertex corresponding to the edge $xy$ in the subdivision graph $S(G)$.
+In $S(G)$, let $V_{1}=V(G)$ be the set of all vertices of $G$, and let $V_{2}=\{w_{xy}:xy\in E(G)\}$ denote the set of all subdivision vertices.
 Lemma 2.1.
@@
 Proof.
-First, we note that $N_{C(G)}(a)\neq N_{C(G)}(b)$ for any $a,b\in V_{2}$.
-Let $x\in V_{2}$.
-If $\psi(x)=y$ for some $y\in V_{2}\backslash\{x\}$, then $\psi(N_{C(G)}(x))=N_{C(G)}(y)$ since automorphisms preserve adjacency.
-Since for every $v\in V_{1}$, $\psi(v)=v$, we have $\psi(N_{C(G)}(x))=N_{C(G)}(x)$.
-Thus, $N_{C(G)}(x)=N_{C(G)}(y)$, which is a contradiction.
+First, we observe that distinct vertices of $V_{2}$ have distinct neighborhoods in $C(G)$.
+Indeed, if $w_{xy},w_{uv}\in V_{2}$ with $w_{xy}\neq w_{uv}$, then $\{x,y\}\neq\{u,v\}$ since $G$ is simple, and hence each edge of $G$ has a unique pair of endpoints.
+Therefore,
+$N_{C(G)}(w_{xy})=\{x,y\}\neq\{u,v\}=N_{C(G)}(w_{uv}).$
+Now, let $x\in V_{2}$.
+If $\psi(x)=y$ for some $y\in V_{2}\setminus\{x\}$, then $\psi(N_{C(G)}(x))=N_{C(G)}(y)$ since automorphisms preserve adjacency.
+Since $N_{C(G)}(x)\subseteq V_{1}$ and $\psi(v)=v$ for every $v\in V_{1}$, it follows that $\psi(N_{C(G)}(x))=N_{C(G)}(x)$.
+Thus, $N_{C(G)}(x)=N_{C(G)}(y)$, a contradiction.
+Since $\psi(u)=u$ for all $u\in V_{1}$, we have $\psi(V_{1})=V_{1}$, and hence $\psi(x)\notin V_{1}$.
+Therefore, $\psi(x)=x$ for all $x\in V_{2}$, and consequently $\psi=id_{C(G)}$.
 ∎
@@
 Proof.
-In $C(G)$, each vertices of $V_{1}$ is of degree $n-1\geq 3$ and each vertices of $V_{2}$ is of degree $2$.
-The rest follows from the fact that $deg_{C(G)}(\phi(v))=deg_{C(G)}(v)$ for any $v\in V(C(G))$.
+In $C(G)$, each vertex of $V_{1}$ has degree $n-1\geq 3$, whereas each vertex of $V_{2}$ has degree $2$.
+Since automorphisms preserve vertex degrees, that is, $deg_{C(G)}(\varphi(v))=deg_{C(G)}(v)$ for any $v\in V(C(G))$, it follows that $\varphi(V_{1})=V_{1}$ and $\varphi(V_{2})=V_{2}$.
 ∎
@@
 Let $\alpha\in Aut(G)$. Define the mapping $\pi_{\alpha}:V(C(G))\rightarrow V(C(G))$ as follows:
-$\pi_{\alpha}(w)=\begin{cases}\alpha(w)&\text{if }w\in V(G)\\ w_{\alpha(x),\alpha(y)}&\text{if }w\in V(C(G))\backslash V(G),w=w_{x,y},x,y\in V(G).\end{cases}$
+$\pi_{\alpha}(w)=\begin{cases}\alpha(w)&\text{if }w\in V(G)\\ w_{\alpha(x)\alpha(y)}&\text{if }w\in V(C(G))\backslash V(G),w=w_{xy},x,y\in V(G).\end{cases}$
 We can see that $\pi_{\alpha}\in Aut(C(G))$.
-If $\{x,y\}\in E(C(G))$ then $\{x,y\}\in E(\overline{G})$ or $\{x,y\}\in E(S(G))$.
-By the definition of $\pi_{\alpha}$, if $\{x,y\}\in E(S(G))$ then $\{\pi_{\alpha}(x),\pi_{\alpha}(y)\}\in E(S(G))$.
-Moreover, we have $\{x,y\}\in E(\overline{G})\iff\{x,y\}\not\in E(G)\iff\{\alpha(x),\alpha(y)\}\not\in E(G)\iff\{\alpha(x),\alpha(y)\}\in E(\overline{G})\iff\{\pi_{\alpha}(x),\pi_{\alpha}(y)\}\in E(\overline{G})$.
+If $xy\in E(C(G))$ then either $xy\in E(\overline{G})$ or $xy\in E(S(G))$.
+By the definition of $\pi_{\alpha}$, if $xy\in E(S(G))$ then $\pi_{\alpha}(x)\pi_{\alpha}(y)\in E(S(G))$.
+Moreover, we have
+$xy\in E(\overline{G})$ $\iff xy\notin E(G)$
+$\iff\alpha(x)\alpha(y)\notin E(G)$
+$\iff\alpha(x)\alpha(y)\in E(\overline{G})$
+$\iff\pi_{\alpha}(x)\pi_{\alpha}(y)\in E(\overline{G}).$
 Thus, if $H=\{\pi_{\alpha}:\alpha\in Aut(G)\}$, then $H\subseteq Aut(C(G))$.
@@
 By Lemma 2.2, we have $f(V_{1})=V_{1}$ since $G$ is a connected graph of order at least $4$.
-If $\alpha=f\restriction V_{1}$, then $\alpha\in Aut(G)$.
+Thus, if $\alpha=f\restriction V_{1}$, then $\alpha\in Aut(G)$.
 Let $\pi_{\alpha}$ be the induced automorphism of $\alpha$ as defined above.
-If we let $l=f^{-1}\pi_{\alpha}$, then $l$ is an automorphism of $C(G)$ such that $l(v)=v$ for every $v\in V_{1}$.
+If we let $l=f^{-1}\circ\pi_{\alpha}$, then $l$ is an automorphism of $C(G)$ such that $l(v)=v$ for every $v\in V_{1}$.
 By Lemma 2.1, $l=id_{C(G)}$.
@@
 Definition 2.4.
-Let $G=(V(G),E(G))$ be a graph where $V(G)=\{v_{1},...,v_{p}\}$ for some integer $p\geq 1$.
-To $G$, we add $p$ new vertices $\{u_{1},...,u_{p}\}$ and $p$ new edges $\{u_{i},v_{i}\}$.
-The new graph is the endline graph of $G$ denoted by $G^{+}$.
-The edges $\{u_{i},v_{i}\}$ are the endline edges of $G^{+}$.
-Let $U_{1}=V(G)$ and $U_{2}=V(G^{+})\backslash V(G)$.
-Let $L(G)$ be the line graph of $G$.
-Fact 2.7.
+[…]
+Fact 2.5.
 The following holds:
@@
   3. (3)
-    (Mirafzal; [25, Theorem 4]) If $G\not\cong C_{n}$ for any $n\geq 3$, then $Aut(S(G))\cong Aut(G)$.
+    (Mirafzal; [14, Theorem 4]) If $G\not\cong C_{n}$ for any $n\geq 3$, then $Aut(S(G))\cong Aut(G)$.
 […]
@@
 Let $u_{i}\in U_{2}$.
-If $\psi(u_{i})=u_{j}$ for some $u_{j}\in U_{2}\backslash\{u_{i}\}$, then $\psi(v_{i})=v_{j}$ since the automorphisms preserve adjacency, which is a contradiction to the given assumption.
+Suppose $\psi(u_{i})=u_{j}$ for some $u_{j}\in U_{2}$ with $j\neq i$.
+Since automorphisms preserve adjacency and $u_{i}$ is adjacent only to $v_{i}\in U_{1}$, it follows that $\psi(v_{i})=v_{j}$.
+However, by assumption $\psi(v_{i})=v_{i}$ for all $v_{i}\in U_{1}$, which is a contradiction.
+Thus $\psi(u_{i})=u_{i}$ for all $u_{i}\in U_{2}$, and hence $\psi=id_{G^{+}}$.
 ∎
@@
 Proof.
-In $G^{+}$, each vertex of $U_{1}$ has a degree at least $2$ and each vertex of $U_{2}$ has a degree $1$. ∎
+In $G^{+}$, each vertex of $U_{1}$ has degree at least $2$, while each vertex of $U_{2}$ has degree $1$.
+Since automorphisms preserve vertex degrees, it follows that $\varphi(U_{1})=U_{1}$ and $\varphi(U_{2})=U_{2}$.
+∎
 Lemma 2.8.
@@
 Since $G^{+}\not\in\{Q,L(Q),P_{2}\}$ as $G^{+}$ has $n$ vertices of degree one (see Fig.
-1), $Aut(M(G))\cong Aut(L(G^{+}))\cong Aut(G^{+})$ by Fact 2.7.
+1), $Aut(M(G))\cong Aut(L(G^{+}))$ by the Fact 2.5(1), and $Aut(L(G^{+}))\cong Aut(G^{+})$ by the Fact 2.5(2).
 Let $\alpha\in Aut(G)$.
@@
 $\pi_{\alpha}(w)=\begin{cases}\alpha(v_{i})&\text{if }w=v_{i}\in U_{1}\\ u_{j}&\text{if }w=u_{i}\in U_{2},v_{j}=\alpha(v_{i}).\end{cases}$
-Then $\pi_{\alpha}\in Aut(G^{+})$, where $\pi_{\alpha}(U_{i})=U_{i}$ for $i\in\{1,2\}$. The rest follows by applying Lemmas 2.8, 2.9, and the arguments in the proof of Lemma 2.5. ∎
-By Lemmas 2.5, 2.10, and Fact 2.7(2,3), we have the following result.
+Then $\pi_{\alpha}\in Aut(G^{+})$, where $\pi_{\alpha}(U_{i})=U_{i}$ for $i\in\{1,2\}$. The rest follows by applying Lemmas 2.6, 2.7, and the arguments in the proof of Lemma 2.3. ∎
+By Lemmas 2.3 and 2.8, together with Fact 2.5(2,3), we obtain the following result.
 Theorem 2.9.
 If $G$ is a connected graph of order $n\geq 5$ and $G\not\cong C_{n}$, then $Aut(G)\cong Aut(L(G))\cong Aut(S(G))\cong Aut(C(G))\cong Aut(M(G))$.
+Proof.
+Since $n\geq 5$, we have $G\not\cong P_{2},Q,$ or $L(Q)$.
+Hence, by Fact 2.5(2), $Aut(G)\cong Aut(L(G)).$ Since $G\not\cong C_{n}$, Fact 2.5(3) implies that $Aut(S(G))\cong Aut(G).$ Moreover, by Lemma 2.3 and Lemma 2.8, we have $Aut(C(G))\cong Aut(G)\text{ and }Aut(M(G))\cong Aut(G).$
 3. Distinguishing number and distinguishing index
@@
   1. (1)
-    (Kalinowski, Pilśniak, and Woźniak; [15, Theorem 2.2]) If $n\geq 3$, then $D^{\prime\prime}(G)\leq\lceil\sqrt{\Delta(G)}\rceil$.
+    (Kalinowski, Pilśniak, and Woźniak; [9, Theorem 2.2]) If $n\geq 3$, then $D^{\prime\prime}(G)\leq\lceil\sqrt{\Delta(G)}\rceil$.
   2. (2)
-    (Alikhani and Soltani; [4, Theorems 3.8 and 4.3]) If $n\geq 3$, then $D(S(G))\leq\lceil\sqrt{\Delta(G)}\rceil$ and $D^{\prime}(S(G))\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$.
+    (Alikhani and Soltani; [3, Theorems 3.8 and 4.3]) If $n\geq 3$, then $D(S(G))\leq\lceil\sqrt{\Delta(G)}\rceil$ and $D^{\prime}(S(G))\leq min\{r:r^{2}-r\geq D^{\prime}(G)\}$.
   3. (3)
-    (Kalinowski and Pilśniak; [16, Propositions 5 and 6]) $D^{\prime}(C_{t})=2$ and $D^{\prime}(K_{t})=2$ for $t\geq 6$.
+    (Kalinowski and Pilśniak; [10, Propositions 5 and 6]) $D^{\prime}(C_{t})=2$ and $D^{\prime}(K_{t})=2$ for $t\geq 6$.
   4. (4)
-    (Alikhani and Soltani; [3, Theorem 2.3]) If $G\not\in\{P_{2},L(Q)\}$, then $D(L(G))=D^{\prime}(G)$ (see Fig. 1).
+    (Alikhani and Soltani; [2, Theorem 2.3]) If $G\not\in\{P_{2},L(Q)\}$, then $D(L(G))=D^{\prime}(G)$ (see Fig. 1).
   5. (5)
-    (Kalinowski and Pilśniak; [16]) If $n\geq 3$, then $D^{\prime}(G)\leq\Delta(G)$ unless $G$ is $C_{3},C_{4},$ or $C_{5}$.
+    (Kalinowski and Pilśniak; [10]) If $n\geq 3$, then $D^{\prime}(G)\leq\Delta(G)$ unless $G$ is isomorphic to $C_{3},C_{4},$ or $C_{5}$.
+Proposition 3.2.
+For any integer $n\geq 4$, we have $D^{\prime}(C(K_{n}))=2$.
+Proof.
+By Lemma 2.3, we have $Aut(C(K_{n}))\cong Aut(K_{n})\cong S_{n}$.
+Thus, $|Aut(C(K_{n}))|=n!>1$ and so $C(K_{n})$ is not asymmetric.
+Consequently, $D^{\prime}(C(K_{n}))>1$ and it suffices to construct an edge coloring $c:E(C(K_{n}))\to\{1,2\}$ such that the only automorphism of $C(K_{n})$ that preserves $c$ is $id_{C(K_{n})}$.
+Let $V(K_{n})=\{v_{1},\dots,v_{n}\}$.
+Recall that for each edge $v_{i}v_{j}$ of $K_{n}$, $w_{v_{i}v_{j}}$ is the subdivision vertex introduced on that edge.
+Define the edge coloring $c:E(C(K_{n}))\to\{1,2\}$ such that for each $v_{i}\in V_{1}$, the incident edges $v_{i}w_{v_{i}v_{j}}$ are colored $2$ if $j<i$ and $1$ if $j>i$ (see Figure 2).
+[…]
+Figure 2. A schematic representation of the edge coloring of $C(K_{n})$. Edges $v_{i}w_{v_{i}v_{j}}$ are colored 22 (blue) if $j<i$ and 11 (red) if $j>i$.
+Let $\varphi\in\mathrm{Aut}(C(K_{n}))$ be an automorphism that preserves $c$.
+For each $i\in\{1,...,n\}$, let $d_{2}(v_{i})$ denote the number of edges incident to $v_{i}$ that are colored with $2$.
+By construction, $d_{2}(v_{i})=i-1\text{ for all }i\in\{1,\dots,n\}.$ Hence, the values $d_{2}(v_{i})$ are distinct.
+Since $\varphi$ preserves $c$, we have $\varphi(v_{i})=v_{i}$ for all $v_{i}\in V(K_{n})$.
+Thus, by Lemma 2.1, $\varphi=id_{C(K_{n})}$.
+∎
+Proposition 3.3.
+For any integer $n\geq 4$, we have $D^{\prime}(C(C_{n}))=2$.
+Proof.
+Since $Aut(C(C_{n}))\cong Aut(C_{n})\cong D_{n}$ by Lemma 2.3 where $D_{n}$ is the Dihedral group, the graph $C(C_{n})$ is not asymmetric, and thus $D^{\prime}(C(C_{n}))>1$.
+We show $D^{\prime}(C(C_{n}))\leq 2$.
+Let $V(C(G))=V_{1}\cup V_{2}$, where $V_{1}=V(C_{n})$ and $V_{2}=\{w_{xy}:xy\in E(C_{n})\}$ be the set of subdivision vertices.
+Since $n\geq 4$, we can select distinct $v_{1},v_{2},v_{3}\in V_{1}$ so that $v_{1}v_{2},v_{2}v_{3}\in E(C_{n})$.
+Define an edge coloring $c:E(C(C_{n}))\to\{0,1\}$ as follows (see Figure 3):
+  1. (1)
+    $c(v_{1}w_{v_{1}v_{2}})=1$ and $c(w_{v_{1}v_{2}}v_{2})=1$.
+  2. (2)
+    $c(v_{3}w_{v_{2}v_{3}})=1$.
+  3. (3)
+    $c(e)=0$ otherwise.
+Let $\phi\in Aut(C(C_{n}))$ be an automorphism that preserves $c$. Let $d_{1}(u)$ denote the number of edges incident to $u$ colored with $1$. We can observe the following:
+  • The vertices $w_{v_{1}v_{2}}$ and $w_{v_{2}v_{3}}$ are the unique vertices with $d_{1}(w_{v_{1}v_{2}})=2$ and $d_{1}(w_{v_{2}v_{3}})=1$.
+So $\phi(w_{v_{1}v_{2}})=w_{v_{1}v_{2}}$ and $\phi(w_{v_{2}v_{3}})=w_{v_{2}v_{3}}$.
+Since automorphism preserves neighbors, we have $\phi(\{v_{1},v_{2}\})=\{v_{1},v_{2}\}$ and $\phi(\{v_{2},v_{3}\})=\{v_{2},v_{3}\}$.
+  • Among the neighbors $v_{2}$ and $v_{3}$ of $w_{v_{2}v_{3}}$, $v_{3}$ is the only vertex with $d_{1}(v_{3})=1$.111note that $d_{1}(v_{2})$ also is $1$, but $v_{2}\in N(w_{v_{1}v_{2}})$ Thus $\phi(v_{3})=v_{3}$.
+Consequently, $\phi(v_{2})=v_{2}$.
+Since $\phi(\{v_{1},v_{2}\})=\{v_{1},v_{2}\}$, we have $\phi(v_{1})=v_{1}$.
+Since $\phi$ fixes three consecutive vertices $v_{1},v_{2},$ and $v_{3}$ of the cycle $C_{n}$, $\phi$ must fix all vertices of $C_{n}$. By Lemma 2.1, $\phi=id_{C(C_{n})}$. ∎
+[…]
+Figure 3.
+A schematic representation of the edge coloring of $C(C_{n})$.
+The edges $v_{1}w_{v_{1}v_{2}},w_{v_{1}v_{2}}v_{2}$, and $v_{3}w_{v_{2}v_{3}}$ are colored with 11 (red), and the rest of the edges are colored with 0 (black).
 Theorem 3.4.
@@
 Proof.
-Since $G$ is connected, we have $\Delta(G)\geq 2$.
-Hence, we have at least $\lceil\sqrt{2}\rceil=2$ colors available to color $C(G)$.
+Since $G$ is connected, we have $\Delta(G)\geq 2$. Hence,
+$\left\lceil\sqrt{\Delta(G)}\right\rceil\geq\left\lceil\sqrt{2}\right\rceil=2.$
 For $k\geq 0$, let $S_{k}(x)$ denote a sphere of radius $k$ with a center $x$ in $G$, that is, the set of all vertices at distance $k$ from $x$.
 Case (i): $G=K_{n}$ or $C_{n}$.
-If $n\geq 6$, then $D^{\prime}(G)=2\leq\lceil\sqrt{\Delta(G)}\rceil$ by Fact 3.1(3).
-We assume $n$ is either $4$ or $5$.
-In both cases, we can color the edges of $C(G)$ with two available colors so that the edge coloring is preserved only by $id_{G}$ (see Fig.
-2).
-$C(C_{4})$$C(C_{5})$$C(K_{4})$$C(K_{5})$
-Figure 2. Edge colorings of $G$ with $2$ colors that are only preserved by $id_{G}$ where $G\in\{C(C_{4}),C(C_{5}),C(K_{4}),C(K_{5})\}$.
+By Propositions 3.2 and 3.3, $D^{\prime}(C(G))=2$. Since $\left\lceil\sqrt{\Delta(G)}\right\rceil\geq 2,$ we obtain $D^{\prime}(C(G))\leq\left\lceil\sqrt{\Delta(G)}\right\rceil$.
 Case (ii): $G$ contains a cycle but $G\not\in\{C_{n},K_{n}\}$.
 Thus, we can choose a vertex $v_{0}$ lying on a cycle such that the sphere $S_{2}(v_{0})$ is non-empty. Consider a BFS-spanning tree $T$ of $G$ rooted at $v_{0}$. For a vertex $v\in T$, we denote
-$X_{v,u}=(\{v,w_{v,u}\},\{w_{v,u},u\})$ if $w_{v,u}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)$
+$X_{v,u}=(\{v,w_{vu}\},\{w_{vu},u\})$ if $w_{vu}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)$
 and
-$N(v)=\{X_{v,u}:w_{v,u}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)\}$.
+$N(v)=\{X_{v,u}:w_{vu}\in V_{2}\text{ subdivides the edge }\{v,u\}\in E(T)\}$.
 Step 1: We color all edges of the subdivision graph of $T$ and certain edges of $C(G)$.
   • We pick $v_{1},v_{2}\in S_{1}(v_{0})$ and assign the pair $(1,1)$ of colors to $X_{v_{0},v_{1}}$ and $X_{v_{0},v_{2}}$.
+Without loss of generality, we choose $v_{1}\in S_{1}(v_{0})$ such that $v_{1}$ has a neighbor in $S_{2}(v_{0})$.
 Since the degree of $v_{0}$ is at most $\Delta(G)$ and the pair $(1,1)$ is used twice to color the elements of $N(v_{0})$, we can color every pair of $N(v_{0})\setminus\{X_{v_{0},v_{1}},X_{v_{0},v_{2}}\}$ with different pairs of colors from $(1,1)$.
-We modify an algorithm due to Kalinowski–Pilśniak–Woźniak [15] to color all the edges of the subdivision graph of $T$ so that $v_{0}$ is the only vertex such that the pair $(1,1)$ appears twice in $N(v_{0})$.
+Inspired by an algorithm due to Kalinowski–Pilśniak–Woźniak [9, Proof of Theorem 2.2], we modify the algorithm to color all the edges of the subdivision graph of $T$ so that $v_{0}$ is the only vertex such that the pair $(1,1)$ appears twice in $N(v_{0})$.
 Hence $v_{0}$ will be fixed by every color-preserving automorphism, and all vertices in $N(v_{0})$ will be fixed, except possibly $v_{1}$ and $v_{2}$.
@@
     $T_{v_{i}}=\{X_{v_{i},u}\in N(v_{i}):u\in S_{2}(v_{0})\}.$
+    Note that $T_{v_{1}}\neq\emptyset$ since $v_{1}$ has a neighbor in $S_{2}(v_{0})$, while $T_{v_{2}}$ may possibly be empty.
     In order to distinguish $v_{1}$ and $v_{2}$, we need to take the following steps:
@@
       We color the elements of $T_{v_{1}}$ and $T_{v_{2}}$ with the same pairs so that none of them contains the pair $(1,1)$.
-[…]
-      If $X_{v_{1},u}\in T_{v_{1}}$, then $u\in S_{2}(v_{0})$.
-Thus, $\{v_{0},u\}\in E(\overline{G})\subseteq E(C(G))$.
-We color all such edges with 1.
-On the other hand, we color all edges of the set $\{\{v_{0},u\}\in E(C(G)):X_{v_{2},u}\in T_{v_{2}}\}$ with 00.
+    • ($\ast\ast$)
+      If $X_{v_{1},u}\in T_{v_{1}}$, then $u\in S_{2}(v_{0})$. Hence $v_{0}u\in E(\overline{G})\subseteq E(C(G))$. Let
+      $Z_{v_{1}}=\{v_{0}u\in E(C(G)):X_{v_{1},u}\in T_{v_{1}}\}.$
+      We color all edges of $Z_{v_{1}}$ with $1$.
+      Similarly, we color all edges of the set
+      $Z_{v_{2}}=\{v_{0}u\in E(C(G)):X_{v_{2},u}\in T_{v_{2}}\}$
+      with 0 (if this set is nonempty). In both cases (i.e., $Z_{v_{2}}=\emptyset$ or $Z_{v_{2}}\neq\emptyset$), the resulting color pattern in $C(G)$ distinguishes $v_{1}$ and $v_{2}$.
 Claim 3.5.
@@
 Proof.
-[…]
+Otherwise, $\phi(v_{1})=v_{2}$.
+Therefore, if $u\in S_{2}(v_{0})\cap N_{T}(v_{1})$, where $N_{T}(v)$ is the open neighborhood of a vertex $v$ in the tree $T$, then $\phi(u)\in S_{2}(v_{0})\cap N_{T}(v_{2})$.
+Since $uv_{0}\in E(C(G))$ and $f(uv_{0})=1$, we have $\phi(u)\phi(v_{0})\in E(C(G))$ and $f(\phi(u)\phi(v_{0}))=f(\phi(u)v_{0})=1$.
+This contradicts $(\ast\ast)$.
+∎
   • Suppose that for some integer $k$, every element $u\in\bigcup_{i\leq k}S_{i}(v_{0})$ is colored so that it is fixed by every color-preserving automorphism.
@@
 Continuing in this fashion, we color all the edges of $S(T)$.
-Step 2: We color all the remaining edges in $C(G)$ with 00.
-Step 3: Using Lemmas 2.3 and 2.4, we prove that the coloring that we defined above (say $c$) is preserved only by $id_{C(G)}$.
-  • Let $\psi\in Aut(C(G))$ so that $\psi$ preserves $c$.
+Step 2: We color all the remaining edges in $C(G)$ with 0.
+Step 3: Using Lemmas 2.1 and 2.2, we prove that the coloring we defined above (say, $c$) is preserved only by $id_{C(G)}$.
+  • Let $\psi\in Aut(C(G))$ be an automorphism that preserves $c$.
 By Lemma 2.2, $\psi(V_{i})=V_{i}$ for $i\in\{1,2\}$.
-Since all vertices of $T$ are fixed by $\psi$, we have $\psi(u)=u$ for all $u\in V_{1}$.
-By Lemma 2.3, $\psi$ is $id_{C(G)}$.
+Since all vertices of $T$ are fixed by $\psi$ and the vertices of $T$ correspond to $V_{1}$, it follows that $\psi(u)=u$ for every $u\in V_{1}$.
+By Lemma 2.1, we conclude that $\psi=id_{C(G)}$.
 Case (iii): $G$ is a tree.
@@
   1.
-Case A: If $G$ has a central vertex $v_{0}$, then every element of $N(v_{0})$ obtains a different pair of colors (where $N(v)$ is defined in $Case(ii)$ for any vertex $v$ in the tree $G$).
-This is possible as we have at least $\Delta(G)$ different pairs of colors available.
-Hence, all vertices adjacent to $v_{0}$ in $G$ are fixed by every automorphism of $G$, preserving this coloring.
-If $v\in S_{k}(v_{0})$, then all the elements of $N(v)\backslash\{X_{v,u}:u\in S_{k-1}(v_{0})\}$ obtain a different pair of colors.
-By induction on the distance from $v_{0}$, all vertices of $G$ are fixed.
+Case A: Suppose $G$ has a central vertex $v_{0}$.
+If $S_{2}(v_{0})\neq\emptyset$, we consider $G$ rooted in $v_{0}$ and apply the coloring argument of Case (ii), which fixes vertices successively according to their distance from $v_{0}$.
+If $S_{2}(v_{0})=\emptyset$, then $G\cong K_{1,n-1}$.
+In this case, it is known that
+    $D^{\prime}(C(K_{1,n-1}))=\left\lceil\sqrt{n-1}\right\rceil=\left\lceil\sqrt{\Delta(K_{1,n-1})}\right\rceil.$
+    Thus, the desired bound holds.
   2.
-Case B: If $G$ has a central edge $e_{0}=\{x,y\}$, let $T_{1}$, $T_{2}$ be subtrees obtained by deleting $e_{0}$.
-If we put different colors on the edges $\{x,w_{x,y}\}$ and $\{w_{x,y},y\}$, then $x$ and $y$ are fixed by every automorphism.
+Case B: If $G$ has a central edge $e_{0}=xy$, let $T_{1}$, $T_{2}$ be subtrees obtained by deleting $e_{0}$.
+If we put different colors on the edges $xw_{xy}$ and $w_{xy}y$, then $x$ and $y$ are fixed by every automorphism.
 We color $T_{i}$ using the same method as in Case A for $i\in\{1,2\}$.
+Proposition 3.6.
+For all integers $m,n\geq 3$,
+$D^{\prime}(C(K_{m,n}))\leq\max(D^{\prime}(K_{m}),D^{\prime}(K_{n}))+1.$
+Proof.
+Let $V=\{v_{1},\dots,v_{n}\}$ and $U=\{u_{1},\dots,u_{m}\}$ be the bipartitions of $K_{m,n}$, and let $G=C(K_{m,n})$.
+Then $G[V]\cong K_{n}$ and $G[U]\cong K_{m}$ where $G[V]$ and $G[U]$ are the induced subgraphs of $G$ with respect to $V$ and $U$ respectively.
+Let $G_{n}:=G[V]$ and $G_{m}:=G[U]$ (see Figure 4).
+[…]
+Figure 4. A schematic representation of $C(K_{n,m})$. Representative edges of the cliques $K_{n}$ and $K_{m}$ are drawn, and each edge $v_{i}u_{j}$ is subdivided by a vertex $w_{v_{i}u_{j}}$.
+Suppose that $f_{n}:E(G_{n})\to\{1,\dots,D^{\prime}(K_{n})\}$ and $f_{m}:E(G_{m})\to\{1,\dots,D^{\prime}(K_{m})\}$ are distinguishing edge colorings of $G_{n}$ and $G_{m}$, respectively.
+Let $d=\max(D^{\prime}(K_{m}),D^{\prime}(K_{n}))$.
+Define a coloring $f:E(G)\to\{1,\dots,d+1\}$ as follows:
+[…]
+Claim 3.7.
+$f$ is a distinguishing edge coloring of $G$.
+Proof.
+Let $\phi\in\mathrm{Aut}(G)$ be an automorphism that preserves $f$.
+By Lemma 2.2, $\phi(V\cup U)=V\cup U$ and $\phi(V_{2})=V_{2}$.
+For each $v_{i}\in V$, all edges of the form $v_{i}w_{v_{i}u_{j}}$ are colored $d+1$ with respect to $f$, whereas for each $u_{j}\in U$, no edge of the form $w_{v_{i}u_{j}}u_{j}$ is colored $d+1$.
+Thus, the property
+[…]
+characterizes vertices of $V$.
+Since $\phi$ preserves $f$, it follows that $\phi(V)=V$ and $\phi(U)=U$.
+Define $\psi_{V}:=\phi\restriction V$ and $\psi_{U}:=\phi\restriction U$, i.e., the restrictions of $\phi$ to $V$ and $U$ respectively.
+It is clear that $\psi_{V}\in Aut(G_{n})$ and $\psi_{U}\in Aut(G_{m})$.
+[^2] Let $v_{i}v_{j}\in E(G_{n})$.
+Then $v_{i}v_{j}\in E(G)$ and both endpoints $v_{i}$ and $v_{j}$ lie in $V$.
+Since $\phi\in Aut(G)$, we have $\phi(v_{i})\phi(v_{j})\in E(G)$.
+Since $\phi(V)=V$, it follows that $\phi(v_{i}),\phi(v_{j})\in V$, and hence $\phi(v_{i})\phi(v_{j})\in E(G_{n})$.
+Thus, $\psi_{V}$ preserves adjacency in $G_{n}$, and hence $\psi_{V}\in Aut(G_{n})$.
+Similarly, $\psi_{U}\in Aut(G_{m})$.
+Fix any $v_{i}v_{j}\in E(G_{n})$.
+Then $\psi_{V}(v_{i})\psi_{V}(v_{j})\in E(G_{n})$ where $\psi_{V}(v_{i})$ and $\psi_{V}(v_{j})$ belong to $V$.
+Since $\psi_{V}\in Aut(G_{n})$ and automorphisms preserve adjacency, the endpoints of $\psi_{V}(v_{i}v_{j})$ are $\psi_{V}(v_{i})$ and $\psi_{V}(v_{j})$.
+Now $G_{n}\cong K_{n}$ is a simple graph.
+Thus, we have
+$\psi_{V}(v_{i}v_{j})=\psi_{V}(v_{i})\psi_{V}(v_{j}).$
+Since $\phi$ preserves $f$, we obtain
+$f_{n}(\psi_{V}(v_{i}v_{j}))=f_{n}(\psi_{V}(v_{i})\psi_{V}(v_{j}))=f(\psi_{V}(v_{i}v_{j}))=f(\phi(v_{i}v_{j}))=f(v_{i}v_{j})=f_{n}(v_{i}v_{j}).$
+Thus, $\psi_{V}$ preserves $f_{n}$.
+Similarly, $\psi_{U}$ preserves $f_{m}$.
+We know $f_{n}$ and $f_{m}$ are distinguishing edge colorings of $G_{n}$ and $G_{m}$ respectively.
+Thus, $\psi_{V}$ and $\psi_{U}$ fixes all vertices of $V$ and $U$ respectively.
+Consequently, $\phi$ fixes $V\cup U$ pointwise.
+By Lemma 2.1, we conclude that $\phi=id_{C(K_{m,n})}$.
+∎
 Theorem 3.8.
@@
 Let $f:V(G)\cup E(G)\to\{1,...,D^{\prime\prime}(G)\}$ be a total coloring of $G$ (not necessarily proper) that is preserved only by $id_{G}$.
-We define $\tilde{f}:V(C(G))\to\{1,...,D^{\prime\prime}(G)\}$ so that $\tilde{f}(x)=f(x)$ for all $x\in V_{1}$ and $\tilde{f}(w_{x,y})=f(\{x,y\})$ for all $w_{x,y}\in V_{2}$.
+We define $\tilde{f}:V(C(G))\to\{1,...,D^{\prime\prime}(G)\}$ so that $\tilde{f}(x)=f(x)$ for all $x\in V_{1}$ and $\tilde{f}(w_{xy})=f(xy)$ for all $w_{xy}\in V_{2}$.
 We claim that $\tilde{f}$ is preserved only by $id_{C(G)}$.
@@
   1. (1)
-    if $v\in V_{1}$, then $\tilde{f}(\psi(v))=\tilde{f}(v)\implies f(\psi(v))=f(v)\implies f(\alpha(v))=f(\pi_{\alpha}(v))=f(v)$.
+    If $v\in V_{1}$, then
+    $\tilde{f}(\psi(v))$ $=\tilde{f}(v)$
+    $\implies f(\psi(v))=f(v)$
+    $\implies f(\alpha(v))=f(\pi_{\alpha}(v))=f(v).$
   2. (2)
-    if $w_{x,y}\in V_{2}$, then $\tilde{f}(\psi(w_{x,y}))=\tilde{f}(w_{x,y})\implies\tilde{f}(\pi_{\alpha}(w_{x,y}))=f(\{x,y\})\implies\tilde{f}(w_{\alpha(x),\alpha(y)})=f(\{x,y\})\implies f(\{\alpha(x),\alpha(y)\})=f(\{x,y\})\implies f(\alpha(\{x,y\}))=f(\{x,y\})$.
+    If $w_{xy}\in V_{2}$, then
+    $\tilde{f}(\psi(w_{xy}))$ $=\tilde{f}(w_{xy})$
+    $\implies\tilde{f}(\pi_{\alpha}(w_{xy}))=f(xy)$
+    $\implies\tilde{f}(w_{\alpha(x)\alpha(y)})=f(xy)$
+    $\implies f(\alpha(x)\alpha(y))=f(xy)$
+    $\implies f(\alpha(xy))=f(xy).$
 So, $\alpha=id_{G}$ since $f$ is preserved only by $id_{G}$. Consequently, $\psi=\pi_{\alpha}=id_{C(G)}$. ∎
@@
 Remark 3.9.
-For any integer $n\geq 1$, we have $D^{\prime}(C(K_{1,n}))=D(C(K_{1,n}))=\lceil\sqrt{n}\rceil=\lceil\sqrt{\Delta(K_{1,n})}\rceil$.
-This shows that the bounds mentioned in Theorems 3.2 and 3.4 are sharp (see Fig.
-3).
+For any integer $n\geq 1$, we have
+$D^{\prime}(C(K_{1,n}))=D(C(K_{1,n}))=\lceil\sqrt{n}\rceil=\lceil\sqrt{\Delta(K_{1,n})}\rceil.$
+Thus, there exist graphs for which the bounds mentioned in Theorems 3.4 and 3.8 cannot be improved (see Fig. 5).
 Theorem 3.10.
@@
 Proof.
-(1). By Facts 3.1(4) and 2.7 (1), $D(M(G))=D(L(G^{+}))=D^{\prime}(G^{+})$.
+Let $E$ be the set of all endline edges of $G^{+}$.
+(1). By Facts 3.1(4) and 2.5 (1), $D(M(G))=D(L(G^{+}))=D^{\prime}(G^{+})$.
 Case(i): $G\not\cong C_{n}$ for any $n\geq 3$.
@@
 Then $D^{\prime}(G^{+})\leq\Delta(G)$ by Fact 3.1(5).
-Let $g:E(G)\rightarrow\{1,...,D^{\prime}(G)\}$ be a coloring of $E(G)$ that is preserved only by $id_{G}$ and $E$ be the set of all endline edges of $G^{+}$.
+Let $g:E(G)\rightarrow\{1,...,D^{\prime}(G)\}$ be a coloring of $E(G)$ that is preserved only by $id_{G}$.
 Define the mapping $g^{\prime}:E(G^{+})\rightarrow\{1,...,D^{\prime}(G)\}$ as follows:
@@
 By Lemma 2.6, $\psi=id_{G^{+}}$.
-Case(ii): $G\cong C_{n}$ for some $n\geq 3$. Then $D(M(G))=D^{\prime}(G^{+})=2=\Delta(G)$. This also shows that the bound mentioned in (1) is sharp.
-(2). Since every line graph is claw-free, and the distinguishing index of a connected claw-free graph is at most $3$ (see [29, Theorem 3.4]), we have $D^{\prime}(M(G))=D^{\prime}(L(G^{+}))\leq 3$. ∎
-4. Total distinguishing chromatic number
-Definition 4.1.
-A Latin square of order $k$ is a $k\times k$ array based on the elements $1,2,...,k$ such that each element occurs exactly once in each row and exactly once in each column.
-A Latin square $M=[m_{i,j}]$ of order $k$ is said to be commutative if $m_{i,j}=m_{j,i}$, for $1\leq i,j\leq k$ and $M$ is said to be idempotent if $m_{i,i}=i$, for $1\leq i\leq k$.
-If the rows of $M$ are just cyclic permutations (one shift of the elements to the left) of the previous row, then $M$ is said to be anti-circulant.
-1 5 2 6 3 7 4
-5 2 6 3 7 4 1
-2 6 3 7 4 1 5
-6 3 7 4 1 5 2
-3 7 4 1 5 2 6
-7 4 1 5 2 6 3
-4 1 5 2 6 3 7
-Table 1. An anti-circulant idempotent commutative Latin square of order $7$.
-Definition 4.2.
-A proper edge coloring of a graph $G=(V(G),E(G))$ is a coloring of $E(G)$ so that any two adjacent edges have different colors.
-The chromatic index of $G$, denoted by $\chi^{\prime}(G)$, is the least number of colors needed for a proper edge coloring of $G$.
-Fix an integer $k$.
-We say that $G$ is $k$-edge choosable if for any assignment $L={L(e)}_{e\in E(G)}$ of the lists of available colors to the edges of $G$, there is a proper edge coloring $f$ of $G$ so that $f(e)\in L(e)$ and $|L(e)|=k$ for all $e\in E(G)$.
-The list chromatic index of $G$, denoted by $\chi_{L}^{\prime}(G)$, is the minimum integer $k$ such that $G$ is $k$-edge choosable.
-Fact 4.3.
-The following holds:
-  1. (1)
-    (Galvin; [14]) If $G$ is a bipartite graph, then $\chi^{\prime}(G)=\chi_{L}^{\prime}(G)=\Delta(G)$.
-  2. (2)
-    An idempotent commutative Latin square (ICLS) of order $k$ exists if and only if $k$ is odd.
-  3. (3)
-    If $M(k)=[m_{i,j}]$ is a Latin square where $m_{i,j}\equiv(i+j)k$ (mod $12k-$ 1), $11\leq m_{i,j}\leq 2k-$ 1, for $11\leq i,j\leq 2k-$ 1, then $M(k)$ is an anti-circulant ICLS of order $12k-$ 1.
-  4. (4)
-    (Xie–He;[35, Theorem 1]) If $G$ is a regular graph of even order $n$ and $\Delta(G)\geq\frac{2n}{3}+\frac{23}{6}$, then $G$ has TCC.
-  5. (5)
-    (Kalinowski–Pilśniak–Woźniak;[15]) Every connected graph $G$ satisfies the inequality $\chi^{\prime\prime}_{D}(G)\leq\chi^{\prime\prime}(G)+1$.
-Furthermore, $\chi_{D}^{\prime\prime}(G)=\chi^{\prime\prime}(G)$ if $\chi^{\prime\prime}(G)\geq\Delta(G)+2$.
-4.1. Central graphs
-Proposition 4.4.
-If $G$ is a connected graph, then $\chi^{\prime\prime}_{D}(C(G))$ is $\Delta(C(G))+1$ or $\Delta(C(G))+2$.
+Case(ii): $G\cong C_{n}$ for some $n\geq 3$.
+Claim 3.11.
+$D^{\prime}(G^{+})=2$.
 Proof.
-It is evident that $\Delta(G)+1\leq\chi^{\prime\prime}(G)\leq\chi^{\prime\prime}_{D}(G)$ for any graph $G$. By Theorem 1.3 and Fact 4.3(5), we have $\chi^{\prime\prime}_{D}(C(G))\leq\Delta(C(G))+2$. ∎
-Theorem 4.5.
-Let $G$ be a connected regular graph of order $n\geq 5$. The following holds:
-  1. (1)
-    If $n$ is odd, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
-  2. (2)
-    If $n$ is even, $G\not\cong K_{n}$, and $\Delta(G)\leq\frac{2n-29}{6}$, then $\chi^{\prime\prime}_{D}(C(G))=\Delta(C(G))+1$.
-Consequently, $C(G)$ is Type 1 if $G$ satisfies the requirement of (1) or (2).
-Proof.
-We note that $\Delta(C(G))=n-1$. We recall that $V_{1}=V(G)$ and $V_{2}=V(C(G))\backslash V(G)$.
-[…]
-Claim 4.6.
-There is a total coloring $g$ of $C(G)$ using $\Delta(C(G))+1$ colors.
-Proof.
-Let $V_{1}=\{u_{1},...,u_{p}\}$ and $V_{2}=\{v_{1},...,v_{q}\}$.
-First, we define a proper total coloring $\pi$ of $\overline{G}$ from $M(k)$ with $\Delta(C(G))+1$ colors.
-Let $\pi(u_{i})=m_{i,i}$ for all $1\leq i\leq n$ and $\pi(\{u_{i},u_{j}\})=m_{i,j}$ for all $1\leq i,j\leq n,i=j$ be such that $\{u_{i},u_{j}\}\in E(\overline{G})$.
-Next, we use Fact 4.3(1) to define a proper edge coloring of the bipartite graph $B=C(G)\backslash E(\overline{G})$.
-We note that for each vertex $u_{i}\in V_{1}$, there are $\Delta(C(G))+1-(deg_{\overline{G}}(u_{i})+1)=\Delta(C(G))-deg_{\overline{G}}(u_{i})$ colors that are not used from $X(u_{i})=\{m_{i,j}:1\leq j\leq\Delta(C(G))+1\}$, i.e., the $i^{th}$- row of $M(k)$.
-Since $G$ is a regular graph, we have $\Delta(C(G))-deg_{\overline{G}}(u_{i})=\Delta(C(G))-\Delta(\overline{G})=\Delta(B)$.
-Fix a vertex $u_{i}\in V_{1}$ and an edge $\{u_{i},v_{j}\}\in E(B)$.
-Let $V(u_{i})=\{m_{i,j}:\{u_{i},u_{j}\}\in E(\overline{G}),1\leq j\leq\Delta(C(G))+1\}$ and $L(\{u_{i},v_{j}\})=X(u_{i})-V(u_{i})$.
-Assume that $L=L(\{u_{i},v_{j}\})_{\{u_{i},v_{j}\}\in E(B)}$ is an assignment of lists of available colors for the edges of the bipartite graph $B$.
-By Fact 4.3(1), there exists a proper edge coloring $\pi_{1}$ of $B$ such that $\pi_{1}(e)\in L(e)$ for all $e\in E_{B}$.
-We define a total coloring $g$ of $C(G)$ using $\Delta(C(G))+1$ colors as follows:
-  • Let $g(\{u_{i},v_{j}\})=\pi_{1}(\{u_{i},v_{j}\})$ for all $\{u_{i},v_{j}\}\in E(B)$,
-  • Let $g(v)=\pi(v)$ for all $v\in V_{1}$ and $g(e)=\pi(e)$ for all $e\in E(\overline{G})$,
-  • If $v_{k}=w_{u_{i},u_{j}}\in V_{2}$ is a vertex that subdivides the edge $\{u_{i},u_{j}\}\in E(G)$, then since $n\geq 5$, let $g(v_{k})$ be any color from the set $\{m_{i,1},...,m_{i,n}\}\backslash\{g(\{v_{k},u_{i}\}),g(\{v_{k},u_{j}\}),g(u_{i}),g(u_{j})\}$.
-We can see that the coloring $g$ is a proper total coloring of $C(G)$. ∎
-Claim 4.7.
-$g$ is only preserved by $id_{C(G)}$.
-Proof.
-Let $\phi\in Aut(C(G))$ be such that $\phi$ preserves $g$.
-By Lemma 2.4, we have $\phi(V_{1})=V_{1}$ and $\phi(V_{2})=V_{2}$.
-Define $f:=g\restriction(V(\overline{G})\cup E(\overline{G}))$ and $\psi:=\phi\restriction V_{1}$.
-Then $\phi\in Aut(\overline{G})$.
-Since $\phi$ preserves $g$, we have the following:
-  • For all $v\in V_{1}$, $g(\phi(v))=g(v)$. Thus, $f(\phi(v))=g(\phi(v))=g(v)=f(v)$ since $\phi(v),v\in V_{1}$.
-  • If $\{v,v^{\prime}\}\in E(\overline{G})$ where $v,v^{\prime}\in V_{1}$, then $\{\phi(v),\phi(v^{\prime})\}\in E(\overline{G})$ where $\phi(v),\phi(v^{\prime})\in V_{1}$.
-Moreover, if $N_{C(G)}(a)$ is the open neighborhood of $a\in V(C(G))$, then
-    $\phi(\{v,v^{\prime}\})\in N_{C(G)}(\phi(v))\cap N_{C(G)}(\phi(v^{\prime}))$,
-    since automorphisms preserve adjacency. Thus, $\{\phi(v),\phi(v^{\prime})\}=\phi(\{v,v^{\prime}\})$ since $G$ is a simple graph. Thus, we have the following:
-    $g(\phi\{v,v^{\prime}\})=g(\{v,v^{\prime}\})$ $\implies g(\{\phi(v),\phi(v^{\prime})\})=g(\{v,v^{\prime}\})$
-    $\implies f(\{\phi(v),\phi(v^{\prime})\})=f(\{v,v^{\prime}\})$
-    $\implies f(\phi\{v,v^{\prime}\})=f(\{v,v^{\prime}\}).$
-Thus, $\psi$ preserves $f$.
-Since all vertices in $V_{1}$ have different colors with respect to $f$, only the identity automorphism of $\overline{G}$ will be the $f$-preserving automorphism of $\overline{G}$.
-Thus, $\phi(u)=\psi(u)=u$ for all $u\in V_{1}$.
-By Lemma 2.3, we have $\phi=id_{C(G)}$.
-∎
-(2).
-Since $G$ is regular and $G\not\cong K_{n}$, $\overline{G}$ is regular with non-zero degree $\Delta(\overline{G})=n-1-\Delta(G)$.
-By assumption, $\Delta(\overline{G})\geq\frac{2n}{3}+\frac{23}{6}$.
-Thus, by Fact 4.3(4), $\overline{G}$ has TCC.
-Consequently, by Fact 4.3(5), $\chi_{D}^{\prime\prime}(\overline{G})\in\{\Delta(\overline{G})+1,\Delta(\overline{G})+2\}$.
-Since $\chi_{D}^{\prime\prime}(C(G))\geq\Delta(C(G))+1$, it suffices to show the following.
-Claim 4.8.
-$\chi_{D}^{\prime\prime}(C(G))\leq\Delta(C(G))+1$.
-Proof.
-Consider the bipartite subgraph $H=C(G)\backslash E(\overline{G})$ of $G$. We note that $\Delta(G)=\Delta(H)$. There are two cases to be considered.
-Case(i): $\chi_{D}^{\prime\prime}(\overline{G})=\Delta(\overline{G})+1$.
-Let $f_{1}$ be a proper total coloring of $\overline{G}$ with colors $11,...,\Delta(\overline{G})+$ 1 such that $f_{1}$ is preserved only by $id_{\overline{G}}$.
-By Fact 4.3(1), there exists a proper edge coloring $f_{2}$ of $H$ with $\Delta(G)$ new colors, say $\Delta(\overline{G})+2,...,\Delta(\overline{G})+\Delta(G)+1$.
-We recall that $\Delta(\overline{G})+\Delta(G)+1=n$.
-Define a proper total coloring $f_{3}$ of $C(G)$ with $n$ colors letting $f_{3}(v)=f_{1}(v)$ for all $v\in V(\overline{G})$, $f_{3}(e)=f_{1}(e)$ for all $e\in E(\overline{G})$, $f_{3}(e)=f_{2}(e)$ for all $e\in E(H)$, and if $v^{\prime}=w_{u_{i},u_{j}}\in V_{2}$ is a vertex that subdivides the edge $\{u_{i},u_{j}\}\in E(G)$, then let $f_{3}(v^{\prime})$ be any color from the set $\{1,...,n\}\backslash\{f_{3}(\{w_{u_{i},u_{j}},u_{i}\}),f_{3}(\{w_{u_{i},u_{j}},u_{j}\}),f_{3}(u_{i}),f_{3}(u_{j})\}$, since $n\geq 5$.
-Case(ii): $\chi_{D}^{\prime\prime}(\overline{G})=\Delta(\overline{G})+2$.
-By the arguments in Case (i), let $f_{1}$ be a proper total coloring of $\overline{G}$ with colors $21,...,\Delta(\overline{G})+$ 2 such that $f_{1}$ is preserved only by $id_{\overline{G}}$ and $f_{2}$ be a proper edge coloring of $H$ with $\Delta(G)$ new colors, say $\Delta(\overline{G})+3,...,\Delta(\overline{G})+2+\Delta(G)$.
-Now, $\Delta(\overline{G})+2+\Delta(G)=n+1$.
-Define a proper total coloring $f_{3}$ of $C(G)$ with $n$ colors as follows:
-  • $f_{3}(v)=f_{1}(v)$ for all $v\in V(\overline{G})$, and $f_{3}(e)=f_{1}(e)$ for all $e\in E(\overline{G})$,
-  • $f_{3}(e)=f_{2}(e)$ for all $e\in E(H)$ such that $f_{2}(e)\neq n+1$,
-  • if $e=\{u,v\}$ such that $u\in V_{1}$ and $v\in V_{2}$, and $f_{2}(e)=n+1$, then let $f_{3}(e)$ be any color from the set $\{1,...,\Delta(\overline{G})+2\}\backslash\{f_{3}(u)\}\cup\{f_{3}(e^{\prime}):e^{\prime}\in E(\overline{G}),u\text{ is incident to }e^{\prime}\}$.
-This is possible since $deg_{\overline{G}}(u)\leq\Delta(\overline{G})$.
-  • If $v^{\prime}=w_{u_{i},u_{j}}\in V_{2}$ subdivides the edge $\{u_{i},u_{j}\}\in E(G)$, then we define $f_{3}(v^{\prime})$ as in Case (i).
-Since $f_{1}$ is a proper total coloring of $\overline{G}$ that is preserved only by $id_{\overline{G}}$, we have that $f_{3}$ is only preserved by $id_{C(G)}$ following the arguments of Claim 4.7.
-Consequently, $\chi_{D}^{\prime\prime}(C(G))\leq n=\Delta(C(G))+1$.
-∎
-This concludes the proof of Theorem 4.5(2). ∎
-4.2. Subdivision graphs
-Proposition 4.9.
-Let $G$ be a connected graph of order $n\geq 5$.
-Then $\chi^{\prime\prime}_{D}(S(G))\in\{\Delta(S(G))+1,\Delta(S(G))+2\}$.
-Moreover, if there exists a vertex $x$ of $G$ for which $\phi(x)=x$ for all $\phi\in Aut(G)$, then $\chi^{\prime\prime}_{D}(S(G))$ is $\Delta(S(G))+1$.
-Proof.
-Since $S(G)$ is a bipartite graph with bipartitions $V_{1}$ and $V_{2}$, $S(G)$ has an edge coloring $g$ with $\Delta(S(G))$ colors according to Fact 4.3 (1).
-For all $v\in V_{1}$, since $deg_{S(G)}(v)\leq\Delta(S(G))$, we have
-$S_{v}=\{1,...,\Delta(S(G))+1\}\backslash\{g(e):v\text{ is incident to }e\}\neq\emptyset$.
-Define the total coloring $g_{1}$ of $S(G)$ with $\Delta(S(G))+1$ colors letting $g_{1}(e)=g(e)$ for all $e\in E(S(G))$, $g_{1}(v)$ be any color from $S_{v}$ for all $v\in V_{1}$, and $g_{1}(w)$ be any color out of $\Delta(S(G))+1$ colors different from the color of its neighbors and edges incident on it for all $w\in V_{2}$.
-Suppose that there exists a vertex $x$ of $G$ for which $\phi(x)=x$ for all $\phi\in Aut(G)$.
-In this case, any neighbor of $x$ has to be mapped into itself because every edge incident to $x$ has distinct color.
-Hence, by induction on the distance from $x$, we can conclude that $g_{1}$ is only preserved by $id_{S(G)}$.
-Assume that there are no vertices $x$ of $G$ for which $\phi(x)=x$ for all $\phi\in Aut(G)$.
-Pick any $x\in V(G)$ and redefine $g_{1}(x)=\Delta(S(G))+2$.
-By the arguments in Case(i), $g_{1}$ is only preserved by $id_{S(G)}$.
-∎
-5. AVD-total chromatic number
-5.1. Central graphs of regular graphs
-Theorem 5.1.
-If $G$ is a regular connected graph with order $n\geq 5$ then the following holds:
-  1. (1)
-    If $n$ is even then $C(G)$ is an AVD Type 2 graph.
-  2. (2)
-    If $n$ is odd, then $C(G)$ satisfies AVD-TCC.
-Proof.
-We note that $\Delta(C(G))=n-1$.
-[…]
-($\ast$) the color class $C_{C(G)}(u_{i})$ contains all colors of $\{m_{i,j}:1\leq j\leq\Delta(C(G))+1\}$ and $m_{i,\Delta(C(G))+2}$ is different for each $i$, $1\leq i\leq n$.
-Since $M(k)$ is anti-circulant (see Def.
-4.1) and ()($\ast$)( ∗ ) hold, we have $C_{C(G)}(u_{i})\neq C_{C(G)}(u_{j})$ for any two distinct vertices $u_{i}$ and $u_{j}$ in $V_{1}$.
-Since for all $u_{i}\in V_{1}$, $deg_{C(G)}(u_{i})\neq 2$ and for all $v_{j}\in V_{2}$, $deg_{C(G)}(v_{j})=2$, we have $C_{C(G)}(u_{i})\neq C_{C(G)}(v_{j})$ for each $u_{i}\in V_{1}$ and $v_{j}\in V_{2}$.
-Thus, $\pi$ is an AVD-total coloring of $C(G)$.
-Since the vertices of maximum degree are adjacent in $C(G)$, we have $\chi_{a}^{\prime\prime}(C(G))=\Delta(C(G))+2$.
-[…]
-5.2. Subdivision graphs
-Recall that $S(G)$ is the subdivision graph of $G$, $V_{1}=V(G)$, and $V_{2}$ is the set of all subdivided vertices.
-We note that every bipartite graph $G$ has ($\Delta(G)+2$)-AVD total coloring where $\Delta(G)$ colors are used to color the edges of $G$ and two new colors are used to color the vertices of the bipartitions.
-Theorem 5.2.
-([7, Theorem 2.3]) If $G$ is a bipartite graph, then $\chi^{\prime\prime}_{a}(G)\leq\Delta(G)+2$.
-Since $S(G)$ is a bipartite graph, $\chi^{\prime\prime}_{a}(S(G))\leq\Delta(S(G))+2$ by the above result.
-However, we observe that if $\Delta(G)\geq 5$, then $\chi^{\prime\prime}_{a}(S(G))=\Delta(S(G))+1$.
-Theorem 5.3.
-If $G$ is a connected graph and $\Delta(G)\geq 5$ then $S(G)$ is $AVD$-Type 1.
-Proof.
-Since $\Delta(G)\geq 5$, we have $\Delta(S(G))=\Delta(G)$.
-Since $\Delta(G)+1=\Delta(S(G))+1\leq\chi^{\prime\prime}_{a}(S(G))$, it suffices to show that $\chi^{\prime\prime}_{a}(S(G))\leq\Delta(G)+1$.
-We construct an AVD-total-coloring $f$ of $S(G)$ with colors from the set $S=\{1,...,\Delta(G)+1\}$.
-By Fact 4.3 (1), there is a proper edge coloring $g$ of $S(G)$ with colors from the set $\{1,...,\Delta(G)\}$ as $S(G)$ is a bipartite graph.
-Let $f(e)=g(e)$ for any $e\in E(S(G))$.
-Let $f(v)=\Delta(G)+1$ for all $v\in V_{1}$.
-Next, we color all the vertices of $V_{2}$ so that if $v^{\prime}=w_{v,w}\in V_{2}$ then $C_{S(G)}(v^{\prime})\neq C_{S(G)}(v)$ and $C_{S(G)}(v^{\prime})\neq C_{S(G)}(w)$ where $C_{S(G)}(x)$ is the color class of $x$ in $S(G)$.
-Fix $v^{\prime}=w_{v,w}\in V_{2}$.
-  1. Case A: $deg_{S(G)}(v)\neq 2$ and $deg_{S(G)}(w)\neq 2$. We note that $f(w)=f(v)=\Delta(G)+1$. Let
-    $R_{v^{\prime}}=\{1,...,\Delta(G)+1\}\backslash\{f(v),f(\{v,v^{\prime}\}),f(\{w,v^{\prime}\})\}$.
-  2. Case B: $deg_{S(G)}(v)=deg_{S(G)}(w)=2$. Let $N_{S(G)}(v)=\{w_{v,w},w_{v,w_{1}}\}$, $N_{S(G)}(w)=\{w_{v,w},w_{v,w_{2}}\}$, and
-    $R_{v^{\prime}}=\{1,...,\Delta(G)+1\}\backslash\{f(v),f(\{v,v^{\prime}\}),f(\{w,v^{\prime}\}),f(\{v,w_{v,w_{1}}\}),f(\{w,w_{v,w_{2}}\})\}$.
-  3. Case C: $deg_{S(G)}(v)=2$ and $deg_{S(G)}(w)\neq 2$. Assume $N_{S(G)}(v)$ as in Case B. Let
-    $R_{v^{\prime}}=\{1,...,\Delta(G)+1\}\backslash\{f(v),f(\{v,v^{\prime}\}),f(\{w,v^{\prime}\}),f(\{v,w_{v,w_{1}}\})\}$.
-In each of the above cases, $R_{v^{\prime}}\neq\emptyset$ since $\Delta(G)\geq 5$. Let $f(v^{\prime})$ be any color from $R_{v^{\prime}}$.
-Finally, $f$ is a total coloring of $S(G)$ and since $C_{S(G)}(v)\neq C_{S(G)}(w_{v,w})$ for any fixed edge $\{v,w_{v,w}\}\in E(S(G))$, $f$ is an AVD-total coloring of $S(G)$. ∎
-5.3. Central graphs of join of two graphs
-Fact 5.4.
-The following holds:
-  1. (1)
-    (Zhang et al. [37]) Let $K_{n}$ be a complete graph of order $n$, $n\geq 3$, then
-    $\chi^{\prime\prime}_{a}(K_{n})=\begin{cases}n+1&\text{if }\,n\equiv 0\text{(mod 2)},\\ n+2&\text{if }\,n\equiv 1\text{(mod 2)}.\end{cases}$
-  2. (2)
-    (Panda et al. [28]) If $G\in\{P_{n},C_{n}\}$, then $\chi^{\prime\prime}_{a}(C(G))=n+1$ if $n$ is even and $\chi^{\prime\prime}_{a}(C(G))\leq n+2$ otherwise.
-  3. (3)
-    (Panda et al. [28]) $\chi^{\prime\prime}_{a}(C(K_{1,n}))=n+2$ and $\chi^{\prime\prime}_{a}(C(K_{n}))=n$ for any integer $n\geq 1$.
-Proposition 5.5.
-The following holds:
-  1. (1)
-    If $C(G_{1})$ and $C(G_{2})$ satisfy AVD-TCC, then $C(G_{1}+G_{2})$ satisfies AVD-TCC.
-  2. (2)
-    The graph $C(K_{m,n})$ satisfies AVD-TCC for any $m,n\geq 2$.
-  3. (3)
-    If $G_{1},G_{2}\in\{P_{n},C_{n},K_{n}\}$, then $C(G_{1}+G_{2})$ satisfies AVD-TCC for any integer $n\geq 2$.
-  4. (4)
-    If $G_{1}$ and $G_{2}$ have the same order, then $C(G_{1}+G_{2})$ satisfies AVD-TCC.
-Proof.
-The proofs are straightforward using some coloring arguments and applying the results mentioned in Fact 5.4, and thus we provided the details in the Appendix. ∎
-6. Total dominator chromatic number
-Theorem 6.1.
-Let $G$ be a connected graph of order $n\geq 5$. The following holds:
-  1. (1)
-    If $\Delta(G)\leq n-3$, then $\chi^{t}_{d}(\overline{G})\leq\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
-  2. (2)
-    If $G$ is a tree, then $\chi^{t}_{d}(C(G))\leq\Delta(C(G))+1$.
-Moreover, the upper bounds of $\chi^{t}_{d}(C(G))$ mentioned in (1) and (2) are sharp.
-Proof.
-(1).
-Let $V_{1}=V(G)=\{v_{1},...,v_{n}\}$.
-Let $V_{2}=V(C(G))\backslash V(G)=\{w_{i,j}:\{v_{i},v_{j}\}\in E(G)\}$ be the set of subdivided vertices and $f=(X_{1},...,X_{l})$ be a TDC (i.e., total dominator coloring) of $C(G)$.
-Then, for every $v_{i}\in V_{1}$, there exists $k\leq l$ such that $X_{k}\subseteq N_{C(G)}(v_{i})$.
-Define $g=(W_{1},...,W_{l})$ where $W_{i}=X_{i}\backslash V_{2}$ for all $1\leq i\leq l$.
-Clearly, $g$ is a proper vertex coloring of $\overline{G}$.
-Case (i): If for every $v_{i}\in V_{1}$, there exists $k\leq l$ such that $X_{k}\subseteq N_{C(G)}(v_{i})$ and $W_{k}=X_{k}\cap V_{1}\neq\emptyset$, then $W_{k}\subseteq N_{\overline{G}}(v_{i})$, and so $g$ is a TDC of $\overline{G}$.
-[…]
-Figure 4. In the figure above, shaded regions contain vertices belonging to $V_{2}$.
-Case (ii): Assume that there exists $v_{i}\in V_{1}$ such that $\bigcup_{1\leq t\leq r}X_{k_{t}}\subseteq N_{C(G)}(v_{i})\cap V_{2}$ for some $k_{1},...,k_{r}\leq l$.
-Since $\Delta(G)<n-2$, this implies $N_{C(G)}(v_{i})\cap V_{1}\neq\emptyset$.
-Let $N_{C(G)}(v_{i})\cap V_{1}=\{m_{1},...,m_{p}\}$ for some positive integer $p$.
-For every $1\leq j\leq p$, assume that $m_{j}\in X_{t_{j}}$ for some $t_{j}\leq l$.
-If for some $1\leq j\leq p$, we have $X_{t_{j}}\cap V_{1}\subseteq N_{C(G)}(v_{i})$, then we are done.
-Otherwise, assume that for all $1\leq j\leq p$, we have $X_{t_{j}}\cap V_{1}\not\subseteq N_{C(G)}(v_{i})$.
-Since $m_{j}\in X_{t_{j}}\cap V_{1}\cap N_{C(G)}(v_{i})$, we have $|X_{t_{j}}\cap V_{1}|\geq 2$.
-Specifically, we have $X_{t_{1}}\cap V_{1}\not\subseteq N_{C(G)}(v_{i})$ and $|X_{t_{1}}\cap V_{1}|\geq 2$.
-Note that $X_{k_{1}}\subseteq V_{2}$, and so $W_{k_{1}}=\emptyset$.
-Redefine $W_{k_{1}}=\{m_{1}\}$ and $W_{t_{1}}=X_{t_{1}}\backslash(V_{2}\cup\{m_{1}\})$.
-Then, $W_{k_{1}}\subseteq N_{\overline{G}}(v_{i})$.
-If we do the above procedure on all $v_{i}\in V_{1}$ such that $\bigcup_{1\leq t\leq r}X_{k_{t}}\subseteq N_{C(G)}(v_{i})\cap V_{2}$ for some $k_{1},...,k_{r}\leq l$, then $g=(W_{1},...,W_{l})$ is a TDC of $\overline{G}$.
-Claim 6.2.
-$\chi^{t}_{d}(C(G))\leq n$. Moreover, the bound $\chi^{t}_{d}(C(G))\leq n$ is sharp.
-Proof.
-Let $T_{n}=\{v_{i}:\{v_{i},v_{n}\}\in E(G)\}$.
-Since $G$ is connected, $T_{n}\not=\emptyset$.
-Pick $k\in\{i:v_{i}\in T_{n}\}$.
-Let $X_{i}=\{v_{i}\}$ for all $i\in\{1,...,n-1\}\backslash\{k\}$, $X_{n}=V_{2}$, and $X_{k}=\{v_{k},v_{n}\}$.111i.e., we color all the vertices of $V_{2}$ with $n$, assign the colors $11,...,n-$ 1 to $v_{1},...,v_{n-1}$, and color $v_{n}$ with $k$.
-For the following reasons, $g=(X_{1},...,X_{n})$ is a TDC of $G$:
-  1. (1)
-    $X_{k}\subseteq N_{C(G)}(w_{v_{n},v_{k}})$, and for any $v\in V_{2}\backslash\{w_{v_{n},v_{k}}\}$, there exists $t\in\{1,...,n-1\}$ where $X_{t}\subseteq N_{C(G)}(v)$.
-  2. (2)
-    Fix $x\in V_{1}\backslash\{v_{n},v_{k}\}$.
-    1. (a)
-      If $\{v_{n},v_{k}\}\subseteq N_{C(G)}(x)$, then $X_{k}\subseteq N_{C(G)}(x)$.
-    2. (b)
-      If exactly one of $v_{n}$ and $v_{k}$ is adjacent to $x$ in $C(G)$, then $x$ is adjacent to some vertex $v_{t}$ in $V_{1}\backslash\{v_{n},v_{k}\}$ since $\Delta(G)\leq n-3$.
-Thus, $X_{t}\subseteq N_{C(G)}(x)$.
-    3. (c)
-      If neither $v_{n}$ nor $v_{k}$ is adjacent to $x$ in $C(G)$, then we follow step (b).
-  3. (3)
-    Let $y\in\{v_{n},v_{k}\}$. Following step (b), there exists $t\in\{1,...,n-1\}$ such that $X_{t}\subseteq N_{C(G)}(y)$.
-We show that the bound $\chi^{t}_{d}(C(G))\leq n$ is sharp. Define
-$G=K_{n}\backslash(\{\{v_{i},v_{i+1}\}\in E(K_{n}):1\leq i\leq n-1\}\cup\{\{v_{n},v_{1}\}\})$
-where the vertices $v_{1},...,v_{n}$ of $K_{n}$ are arranged in a clockwise order (see Fig. 5). Then $\Delta(G)=n-3$. It is enough to show that $\chi^{t}_{d}(C(G))\geq n$.
-$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$
-Figure 5. Graph $G$ when $n=6$.
-We note that for any TDC $f=(X_{1},...,X_{l})$ of $C(G)$, |$|\{X_{i}:X_{i}\subseteq V_{1},|X_{i}|=1\}|\geq n-2$. Otherwise, the following cases can occur.
-Case (i): There exist two color classes of two elements, say $X_{1}=\{v_{i},v_{j}\}$ and $X_{2}=\{v_{k},v_{t}\}$.
-Observe that both $v_{i}\in N_{C(G)}(v_{k})\cap N_{C(G)}(v_{t})$ and $v_{j}\in N_{C(G)}(v_{k})\cap N_{C(G)}(v_{t})$ can’t happen together by the definition of $G$, since $n\geq 5$.
-Without loss of generality, assume that $v_{i}\not\in N_{C(G)}(v_{k})$.
-Thus, $v_{i}$ is adjacent to $v_{k}$ in $G$.
-Then $N_{C(G)}(v_{i})\cap N_{C(G)}(v_{k})=\{w_{i,k}\}$ and $X_{m}\not\subseteq N_{C(G)}(w_{i,k})$ for each $1\leq m\leq l$; a contradiction to the fact that $f$ is a TDC of $C(G)$.
-Case (ii): There exists a color class of $m$ elements, say $X_{1}=\{v_{i_{1}},...,v_{i_{m}}\}$, where $m\geq 3$.
-We observe that for any $k,l\in\{i_{1},...i_{m}\}$ where $k\neq l$, we have $\{v_{k},v_{l}\}\in E(G)$.
-Then $N_{C(G)}(v_{l})\cap N_{C(G)}(v_{k})=\{w_{l,k}\}$, $X_{1}\not\subseteq N_{C(G)}(w_{l,k})$ and the contradiction follows from the arguments of Case(i).
-Therefore, |$|\{f(v_{i}):1\leq i\leq n\}|\geq n-1$. Since at least one new color is needed to color some vertices in $C=\{w_{i,j}:\{v_{i},v_{j}\}\in E(G)\}$, we have $l\geq n$. ∎
-(2). In view of (1), it suffices to show that if $\Delta(T)\geq n-3$, then $\chi^{t}_{d}(C(T))\leq n$. The proof is trivial, and hence we sketched the details of the proof in the Appendix. ∎
-7. Open Problems
-The following questions remain open.
-Question 7.1.
-Does AVD-TCC hold for the central graph of any graph?
-Question 7.2.
-Classify all graphs $G$ (other than regular graphs) so that $C(G)$ is a Type 1 graph.
-Funding The research of the first author was supported by the EKÖP-24-4-II-ELTE-996 University Excellence scholarship program of the Ministry of Culture and Innovation from the source of the National Research, Development and Innovation fund.
+Let $U_{1}=V(G)=\{v_{1},\ldots,v_{n}\}$ and $U_{2}=V(G^{+})\setminus V(G)=\{u_{1},\ldots,u_{n}\}$, where $u_{i}$ is the pendant vertex adjacent to $v_{i}$.
+Since the graph $G^{+}$ admits a nontrivial automorphism, any edge coloring with one color is preserved by a non-identity automorphism.
+Hence $D^{\prime}(G^{+})\geq 2$.
+Thus, it is enough to show that $D^{\prime}(G^{+})\leq 2$.
+Define a coloring $c:E(G^{+})\to\{1,2\}$ as follows:
+$c(e)=\begin{cases}2&\text{if }e=v_{1}v_{2}\text{ or }e=v_{1}u_{1},\\ 1&\text{otherwise}.\end{cases}$
+We show that $c$ is a distinguishing edge coloring.
+  • Let $\psi\in\mathrm{Aut}(G^{+})$ be a $c$-preserving automorphism. By Lemma 2.7, $\psi(U_{i})=U_{i}$ for $i=1,2$. Hence $\psi$ maps $E(G)$ to $E(G)$ and the set of endline edges $E$ to itself.
+  • The edge $v_{1}u_{1}$ is the unique endline edge colored $2$.
+Therefore $\psi(v_{1}u_{1})=v_{1}u_{1}$.
+Since $u_{1}$ is the only vertex of degree $1$ incident with this edge, it follows that $\psi(u_{1})=u_{1}$ and consequently $\psi(v_{1})=v_{1}$.
+  • The edge $v_{1}v_{2}$ is colored differently from any other edge in the cycle graph $G$. Thus $\psi(v_{1}v_{2})=v_{1}v_{2}$. Since $\psi(v_{1})=v_{1}$, it follows that $\psi(v_{2})=v_{2}$.
+  • For $v_{p}\in V(G)\setminus\{v_{1},v_{2}\}$, let $P_{v_{1},v_{2},v_{p}}$ denote the unique path in the cycle between $v_{1}$ and $v_{p}$ that contains $v_{2}$.
+Now, $\psi$ fixes all the vertices of $G$ since for any $v_{p},v_{q}\in V(G)\backslash\{v_{1},v_{2}\}$ such that $v_{p}\neq v_{q}$, the length of $P_{v_{1},v_{2},v_{p}}$ is different from the length of $P_{v_{1},v_{2},v_{q}}$.
+  • Thus, $\psi(u)=u$ for every $u\in U_{1}$. By Lemma 2.6, we conclude that $\psi=id_{G^{+}}$.
+Hence, $D^{\prime}(G^{+})\leq 2$. This concludes the proof of the claim. ∎
+Thus, $D(M(G))=D^{\prime}(G^{+})=2=\Delta(G)$. This also shows that the bound mentioned in (1) is sharp.
+(2). Since every line graph is claw-free, and the distinguishing index of a connected claw-free graph is at most $3$ (see [16, Theorem 3.4]), we have $D^{\prime}(M(G))=D^{\prime}(L(G^{+}))\leq 3$. ∎
+Data availability No data are associated with this article.
+Declarations
+Conflict of interest The author declares that there is no conflict of interest.